### v2

# Desingularization of double covers of regular surfaces

Abstract.
Let $Z$ be a noetherian integral excellent regular scheme of dimension 222. Let $Y$ be an integral normal scheme endowed with a finite flat morphism $Y\to Z$ of degree 222. We give a description of Lipman’s desingularization of $Y$ by explicit equations, leading to a desingularization algorithm for $Y$.
2020 Mathematics Subject Classification:
11G20, 14B05, 14J17, 14J20
I would like to thank INRIA Sud-Ouest and the team Canari for their hospitality during the preparation of a part of this work.
1. Introduction
The existence of the desingularization of excellent surfaces (reduced noetherian excellent schemes of dimension 2) is well known from the theoretical point of view. However, in practice, the implementation of an algorithm of desingularization in full generalities is rather complicated, see [3] and [7] for an overview of various approaches. In view of applications in arithmetic geometry, a lot of work also have been done for smooth projective curves $C$ over a discrete valuation field $K$. The aim is then to find a projective regular scheme over the valuation ring $\mathcal{O}_{K}$ of $K$, with generic fiber isomorphic to $C$, see [8], [15] and the references therein, when $C$ is a cyclic cover of $\mathbb{P}^{1}_{K}$, mostly assuming the residue characteristic of $K$ to be prime to the order of the cover. For hypersurfaces in the affine plane $\mathbb{A}^{2}$ over $\mathcal{O}_{K}$, see also [5] with some powerful applications in number theory.
In this work we deal with a special type of surfaces for which Lipman’s method works very well and is completely explicit.
Theorem 1.1 (Lipman, [17]).
Let $Y$ be a noetherian integral normal excellent surface. Consider a sequence of birational morphisms
$\to Y_{n}\to\dots\to Y_{1}\to Y_{0}:=Y$ (1.1)
where each $Y_{i+1}\to Y_{i}$ is the normalization of the blowing-up of $Y_{i}$ along some singular closed (reduced) point. Then the sequence is necessarily finite.
The surfaces we are interested in are double covers of regular surfaces. For simplicity we only work with noetherian excellent schemes (e.g., algebraic varieties over a field, schemes of finite type over Dedekind domains of characteristic 0 or complete discrete valuation rings, see [9], §8.2.3).
  • Definition 1.2
    Let $Z$ be an integral noetherian regular scheme. We define a double cover of $Z$ as a reduced scheme $Y$ endowed with a finite flat morphism of degree 222
    $\psi:Y\to Z.$
    We call $Y$ a normal double cover of $Z$ if moreover $Y$ is integral and normal.
  • Example 1.3
    Let $C$ be a hyperelliptic curve over $\mathbb{Q}$, with the canonical degree 222 morphism $C\to\mathbb{P}^{1}_{\mathbb{Q}}$. Let $Z=\mathbb{P}^{1}_{\mathbb{Z}}$ and let $W$ be the normalization of $Z$ in the function field $K(C)$ of $C$. Then $Z$ is a regular excellent surface, and $W\to Z$ is a normal double cover of $Z$. The surface $W$ is a so called Weierstrass model of $C$. It is singular in general. A resolution of singularities on $W$ gives a projective regular model of $C$ over $\mathbb{Z}$.
  • Example 1.4
    More generally, let $Z$ be an integral noetherian excellent surface. Let $L$ be a (possibly inseparable) quadratic extension of the function field $K(Z)$ of $Z$ and let $Y$ be the normalization of $Z$ in $L$. Then $Y\to Z$ is a normal double cover. Indeed, $Y\to Z$ is finite because $Z$ is excellent. The scheme $Y$ a noetherian integral normal surface, hence Cohen-MaCaulay, hence flat over $Z$ because the latter is regular.
Now come back to the general situation. The key point in our situation is that if we start with a normal double cover $Y\to Z$ of a regular surface, then in any sequence like (1.1), all the $Y_{i}$’s are again normal double covers of regular surfaces. More precisely (Theorem 3.4), we have a canonical commutative diagram
[…]
where the vertical arrows are normal double covers of regular surfaces, and the lower horizontal arrows are blowing-ups of (reduced) closed points $q_{i}\in Z_{i}$. Each morphism $Y_{i+1}\to Y_{i}$ is a normalized blowing-up, meaning that we blow-up the singular point (with the reduced structure) of $Y_{i}$ lying over $q_{i}$, then we take the normalization.
Moreover, we can give explicit equations for the $Y_{i}$’s as follows. First one can cover $Z_{0}$ by small affine open subsets $V$’s over which $Y_{0}$ is defined by a monic polynomial of degree 222 (Lemma 2.2):
$y^{2}+ay+b=0,\quad a,b\in\mathcal{O}_{Z}(V).$ (1.2)
If $Z_{1}\to Z_{0}$ is the blowing-up of a point $q_{0}\in Z_{0}$, and if $\lambda_{q_{0}}(Y_{0})$ is the multiplicity of $Y_{0}$ at the point lying over $q_{0}$ (Definitions 2.1 and 2.3), we set $r=[\lambda_{q_{0}}(Y_{0})/2]$. Let $s,t$ be a system of coordinates of $Z_{0}$ at $q_{0}$. Then $Y_{1}$ is covered by the affine schemes defined respectively by
$y_{1}^{2}+(a/s^{r})y_{1}+(b/s^{2r})=0,\quad y_{1}=y/s^{r}$
and
$z_{1}^{2}+(a/t^{r})z_{1}+(b/t^{2r})=0,\quad z_{1}=y/t^{r}$
(Proposition 3.3). And we continue in the same way for the $Y_{i}$’s.
The computation of the multiplicity $\lambda_{q}(Y)\in\mathbb{N}$ is straightforward when $2\in\mathcal{O}_{Z,q}^{*}$. Indeed we can reduce to $a=0$ and $b(q)=0$. Then $\lambda_{q}(Y)$ is just the order of $b$ in the regular local ring $\mathcal{O}_{Z,q}$ (Proposition 2.6(2)). If 222 is not invertible, there is no canonical choice of the equation (1.2). We give a simple algorithm (Lemma 2.10) to determine $\lambda_{q}(Y)$ in the case.
As a byproduct of Theorem 3.4 we prove the existence of a simultaneous resolution of singularities for covers of degree 222 of excellent normal surfaces (Corollary 3.12).
When 222 is invertible in $\mathcal{O}_{Z}$, the sequence $(Z_{i})_{i\geq 0}$ is classical and consists in making the branch locus of $Y\to Z$ regular. In the case $Z=\mathbb{P}^{1}_{R}$ over a discrete valuation ring $R$, this is essentially equivalent to locate the branch locus of the generic fiber of $Y\to Z$. The language of clusters ([6], §1.1, [15]) is then more explicit and convenient. The novelty here is that our method works in any (mixed or equal) characteristic. It can also be applied to the case when $Z$ is a regular model over $R$ of a smooth curve $Z_{K}$ over $K:=\mathrm{Frac}(R)$ and $Y_{K}\to Z_{K}$ is a double cover of smooth curves over $K$.
Our initial motivation is to give an easily implementable algorithm to finding a desingularization of the surface $W$ in Example 1, in the spirit of Tate’s algorithm for elliptic curves [21]. Indeed, a regular model of $C$ over $\mathbb{Z}$ allows to compute some important arithmetic invariants of $C$ as its Artin conductor, the Tamagawa number of its Jacobian $J(C)$ as well as the (finite part of) $L$-function of $J(C)$. Applications will be given in [13]. Implementation in PARI/gp [18] by B. Allombert and in MAGMA are ongoing.
Now let us describe the content of this paper. In § 2, we introduce our multiplicity $\lambda_{p}(Y)$ for any double cover $Y\to Z$ of regular scheme and for any point $p\in Y$. This invariant already appeared in various works, when $Z$ is a smooth surface over a perfect field, or in the context of Hironaka’s characteristic polyhedra (see e.g. [4], §5). In our setting it is particularly simple to define.
Normalized blowing-ups are defined in § 3 as blowing-ups followed by a normalization. These are the morphisms appearing in the top line of Sequence (1.1). We prove the main theorem 3.4 about the description of a sequence of normalized blowing-ups. The equation of each $Y_{i}$ is given in Proposition 3.3.
In § 4, we give a bound on the sum of the multiplicities running through the closed points of the exceptional locus of the normalized blowing-up $Y_{1}\to Y$ (Theorem 4.6). The relative dualizing sheaves $\omega_{Y_{i}/S}$ are determined in § 5 when $Z$ is flat of finite type over a regular scheme $S$. This will be used later in [13] to determine a canonical volume form on the Néron model of the Jacobian $J(C)$ of a hyperelliptic curve $C$. Finally in § 6 we apply the above results to describe a desingularization algorithm for the double covers of regular excellent surfaces.
The general setting in this work is the following: $Z$ is an integral noetherian excellent regular scheme, $Y\to Z$ is a double cover. No assumption are made on the residue fields of $Z$. Except in § 2, $Z$ will be of dimension 222.
2. The multiplicity $\lambda_{p}(Y)$
The multiplicity of $Y$ at a closed point $p$ is a local invariant depending only on $\mathcal{O}_{Y,y}$. Let $A$ be an integral noetherian regular ring. Let $B=A[y]$ be an $A$-algebra with $y^{2}+ay+b=0$ for some $a,b\in A$. We will define a multiplicity $\lambda_{\mathfrak{m}}(B)\in\mathbb{N}$ for any maximal ideal $\mathfrak{m}$ of $A$. It is a measure of singularity of $B$ at the maximal ideals lying over $\mathfrak{m}$.
2.1. Definition and basic properties for local rings
Let us first recall some elementary facts on regular rings. Let $\mathfrak{m}$ be a maximal ideal of $A$, let $d=\dim A_{\mathfrak{m}}$ and $k=A/\mathfrak{m}$. Consider
$\mathrm{Gr}_{\mathfrak{m}}(A)=\oplus_{\ell\geq 0}\mathfrak{m}^{\ell}/\mathfrak{m}^{\ell+1}.$
It is canonically a homogeneous $k$-algebra. As $A$ is regular, it is isomorphic to $\mathrm{Sym}_{k}(\mathfrak{m}/\mathfrak{m}^{2})$, the polynomial ring in $d$ variables over $k$.
For any element non zero element $a\in A$, $\operatorname{ord}_{\mathfrak{m}}(a)$, or simply $\operatorname{ord}(a)$ if $\mathfrak{m}$ is implicitly given, is the supremum of the integers $n$ such that $a\in\mathfrak{m}^{n}$. It depends only on the ideal $aA$. By convention $\operatorname{ord}(0)=+\infty$. We have $\operatorname{ord}(ab)=\operatorname{ord}(a)+\operatorname{ord}(b)$ for $a,b\in A\setminus\{0\}$. As $\mathrm{Gr}_{\mathfrak{m}}(A)$ is an integral domain, $\operatorname{ord}$ extends to a valuation on $\mathrm{Frac}(A)$.
  • Definition 2.1
    Keep the above notation. If $\operatorname{ord}_{\mathfrak{m}}(a)=\ell\neq+\infty$, the initial form of $a$ is the image of $a$ in $\mathfrak{m}^{\ell}/\mathfrak{m}^{\ell+1}$. The latter can be viewed as the $k$-vector space of homogeneous polynomials in $d$ variables and of degree $\ell$.
Lemma 2.2.
Let $R$ be a noetherian ring. Let $D$ be an $R$-algebra, free of rank 222 as $R$-module. Then there exists $y\in D$ such that $D=R\oplus Ry$, and we have for some $a,b\in R$
$y^{2}+ay+b=0.$
Moreover, any other $z$ such that $D=R[z]$ satisfies $z=uy+v$ with $u\in R^{*}$ and $v\in R$.
Proof.
Locally on $\operatorname{Spec}R$, $R$ is a direct summand in $D$ ([2], II, §3, Proposition 6 with $M=R$ and $N=D$), then apply op. cit., Cor. 1 to Proposition 12 with $M=D$ and $N=R$, which implies that $R$ is a direct summand in $D$. The existence of the equation on $y$ and the statement on $z$ are obvious. ∎
  • Definition 2.3
    Let $A$ be an integral noetherian regular ring. Let $B$ be a reduced $A$-algebra, free of rank 222 as $A$-module. Write $B=A[y]$ with
    $y^{2}+ay+b=0$
    for some $a,b\in A$.
    Fix a maximal ideal $\mathfrak{m}\subset A$. For each generator $y$ as above, denote by
    $\lambda_{\mathfrak{m}}(y)=\min\{2\operatorname{ord}_{\mathfrak{m}}(a),\operatorname{ord}_{\mathfrak{m}}(b)\}.$
    Define
    $\lambda_{\mathfrak{m}}(B)=\sup_{y}\{\lambda_{\mathfrak{m}}(y)\}\in\mathbb{N}\cup\{+\infty\},$
    where the supremum runs through all generators $y$ as above. Note that when an $y$ is given, we only need to compute $\lambda(y+c)$ for all $c\in A$ to determine $\lambda_{\mathfrak{m}}(B)$.
  • Remark 2.4
    The above multiplicity is very likely twice the invariant $\delta(B)\in\mathbb{Q}_{\geq 0}$ considered by [4], §5, in a much more general context involving Hironaka’s characteristic polyhedra. We use the notation $\lambda_{\mathfrak{m}}(B)$ to keep coherent with our previous work [12], Définition 10 (with $r=1$).
The multiplicity $\lambda_{\mathfrak{m}}(B)$ is local, in the sense that it depends only on $A_{\mathfrak{m}}\to B_{\mathfrak{m}}$ as shown in the lemma below.
Lemma 2.5.
Keep the above notation.
  1. (1)
    We have $\operatorname{ord}_{\mathfrak{m}}=\operatorname{ord}_{\mathfrak{m}A_{\mathfrak{m}}}$ on $\mathrm{Frac}(A)$.
  2. (2)
    Consider the $A_{\mathfrak{m}}$-algebra $B_{\mathfrak{m}}$, then
    $\lambda_{\mathfrak{m}}(B)=\lambda_{\mathfrak{m}A_{\mathfrak{m}}}(B_{\mathfrak{m}}).$
Proof.
(1) This is because $\mathfrak{m}^{n}A_{\mathfrak{m}}\cap A=\mathfrak{m}^{n}$ for all $n\geq 0$.
(2) Clearly $\lambda_{\mathfrak{m}}(B)\leq\lambda_{\mathfrak{m}A_{\mathfrak{m}}}(B_{\mathfrak{m}})$. Let $z$ be a generator of $B_{\mathfrak{m}}$ as $A_{\mathfrak{m}}$-algebra. Let $B=A[y]$. Then there exists $u\in A_{\mathfrak{m}}^{*}$, $v\in A_{\mathfrak{m}}$ such that $y=uz+v$. Approximate $v$ by a sequence $v_{n}\in A$: $v=v_{n}+\alpha_{n}$ with $\alpha_{n}\in\mathfrak{m}^{n}A_{\mathfrak{m}}$. Then $B=A[y-v_{n}]$ and we have, if $n$ is big enough,
$\lambda_{\mathfrak{m}A_{\mathfrak{m}}}(z)=\lambda_{\mathfrak{m}A_{\mathfrak{m}}}(uz)=\lambda_{\mathfrak{m}A_{\mathfrak{m}}}(y-v_{n})=\lambda_{\mathfrak{m}}(y-v_{n})\leq\lambda_{\mathfrak{m}}(B).$
This proves the inverse inequality. ∎
Proposition 2.6.
Let $A,\mathfrak{m}$ and $B$ be as in Definition 2.1, let $k=A/\mathfrak{m}$. Let $\Delta_{B/A}=(a^{2}-4b)A$ be the discriminant ideal.
  1. (1)
    We have $\operatorname{ord}_{\mathfrak{m}}(\Delta_{B/A})=0$ if and only if $A\to B$ is étale at $\mathfrak{m}$.
  2. (2)
    We have $\lambda_{\mathfrak{m}}(B)\leq\operatorname{ord}_{\mathfrak{m}}(\Delta_{B/A})$, with equality when $\mathrm{char}(k)\neq 2$.
  3. (3)
    The following properties are equivalent:
    1. (a)
      $\lambda_{\mathfrak{m}}(B)=+\infty$,
    2. (b)
      $\mathrm{char}(A)=2$, and for any generator $y$ we have $a=0$ and $b$ is a square in the $\mathfrak{m}$-adic completion $\hat{A}_{\mathfrak{m}}$ of $A$.
    3. (c)
      $B\otimes_{A}\hat{A}_{\mathfrak{m}}$ is non-reduced.
    In particular, if $A$ is excellent, then $\lambda_{\mathfrak{m}}(B)$ is finite.
Proof.
(1) is clear.
(2) We have
$\operatorname{ord}_{\mathfrak{m}}(\Delta_{B/A})=\operatorname{ord}_{\mathfrak{m}}(a^{2}-4b)\geq\min\{\operatorname{ord}_{\mathfrak{m}}(a^{2}),\operatorname{ord}_{\mathfrak{m}}(b)\}=\lambda_{\mathfrak{m}}(y).$
Thus $\lambda_{\mathfrak{m}}(B)\leq\operatorname{ord}_{\mathfrak{m}}(\Delta_{B/A})$. When $\mathrm{char}(k)\neq 2$, by Lemma 2.5 we can suppose $2\in A^{*}$. Then $B=A[y+a/2]$ and $\operatorname{ord}_{\mathfrak{m}}(\Delta_{B/A})=\lambda_{\mathfrak{m}}(y+a/2)\leq\lambda_{\mathfrak{m}}(B)$ which implies the equality.
(3) We can suppose $A$ is local.
(a) $\Longrightarrow$ (b) Suppose that $\lambda_{\mathfrak{m}}(B)=+\infty$. Then there exists a sequence $(c_{n})_{n}$ in $A$ such that $\lambda_{\mathfrak{m}}(y+c_{n})\geq 2n$. Thus
$a-2c_{n}\in\mathfrak{m}^{n},\quad b-ac_{n}+c_{n}^{2}\in\mathfrak{m}^{2n}.$
If $\mathrm{char}(A)\neq 2$, then $c_{n}$ converges to some $\hat{c}\in\hat{A}_{\mathfrak{m}}$ and we have $a=2\hat{c}$. As $A\to\hat{A}_{\mathfrak{m}}$ is faithfully flat, $a\in A\cap 2\hat{A}_{\mathfrak{m}}=2A$. Thus $\hat{c}=c\in A$. Then $b=c^{2}$ and $(y-c)^{2}=0$, contradiction because $B$ is reduced by hypothesis. Therefore $\mathrm{char}(A)=2$. Then $a\in\cap_{n}\mathfrak{m}^{n}=\{0\}$ and $b+c_{n}^{2}\in\mathfrak{m}^{2n}$. This implies that $c_{n}$ converges to some $\hat{c}\in\hat{A}_{\mathfrak{m}}$ with $\hat{c}^{2}=b$.
(b) $\Longrightarrow$ (c) is clear. Let us prove (c) $\Longrightarrow$ (a). By Lemma 2.7(1) below, there exists $\hat{\alpha}\in\hat{A}_{\mathfrak{m}}$ such that $(y+\hat{\alpha})^{2}=0$. If $(\alpha_{n})_{n}$ is a sequence in $A$ convergent to $\hat{\alpha}$, then $\lambda_{\mathfrak{m}}(y+\alpha_{n})$ tends to +$\infty$+ ∞.
Finally, if $A$ is excellent then so is $B$, hence $B\otimes_{A}\hat{A}_{\mathfrak{m}}$ is reduced. ∎
Lemma 2.7.
Let $R$ be an integral normal domain, let $D=R[T]/(T^{2}+aT+b)$.
  1. (1)
    The ring $D$ is not integral if and only if $T^{2}+aT+b$ is reducible in $R[T]$.
  2. (2)
    The ring $D$ is non-reduced if and only if $T^{2}+aT+b=(T+c)^{2}$ for some $c\in A$.
Proof.
The “if” part in both statements are obvious. Let us prove the converse. Let $K=\mathrm{Frac}(R)$. As $D$ is flat over $R$, $D\to D\otimes_{R}K$ is injective. If $D$ is non-integral, then $D\otimes_{R}K$ is non-integral and we have $T^{2}+aT+b=(T+c_{1})(T+c_{2})$ for some $c_{1},c_{2}\in K$. This implies that the $c_{i}$’s are integral over $R$, hence belong to $R$. The proof is similar for the non-reduced case. ∎
  • Remark 2.8
    Keep the notation of Proposition 2.6. We have $\Delta_{B/A}=0$ if and only if $B$ is integral and $\mathrm{Frac}(B)/\mathrm{Frac}(A)$ is an inseparable extension.
Proposition 2.9.
Keep the notation of Definition 2.1.
  1. (1)
    The Artinian ring $B/\mathfrak{m}B$ is reduced if and only if $\lambda_{\mathfrak{m}}(B)=0$. If $k:=A/\mathfrak{m}$ is perfect or of $\mathrm{char}(k)\neq 2$, then this is equivalent to $A\to B$tale above $\mathfrak{m}$.
  2. (2)
    The ring $B$ is regular at all maximal ideals lying over $\mathfrak{m}$ if and only if $\lambda_{\mathfrak{m}}(B)\leq 1$.
  3. (3)
    If $\lambda_{\mathfrak{m}}(B)\geq 1$, then there is only one maximal ideal $\mathfrak{n}\subset B$ lying over $\mathfrak{m}$, $A/\mathfrak{m}\to B/\mathfrak{n}$ is an isomorphism and we have $B_{\mathfrak{n}}=B_{\mathfrak{m}}$.
Proof.
By Lemma 2.5 we can suppose $A$ is local. Suppose $\mathrm{char}(k)=2$ (otherwise the proof is easier). Let $\lambda_{\mathfrak{m}}(y)=\lambda_{\mathfrak{m}}(B)$.
If $\lambda_{\mathfrak{m}}(B)=0$ and $\operatorname{ord}_{\mathfrak{m}}(a)=0$, then $A\to B$ is étale and $B$ is regular. Suppose $\lambda_{\mathfrak{m}}(B)=0$ and $\operatorname{ord}_{\mathfrak{m}}(a)>0$. Then $B/\mathfrak{m}B\simeq k[T]/(T^{2}+\bar{b})$. If $T^{2}+\bar{b}\in k[T]$ is reducible, there exists $\alpha\in A^{*}$ such that $b-\alpha^{2}\in\mathfrak{m}$. This implies that $\lambda_{\mathfrak{m}}(B)\geq\lambda_{\mathfrak{m}}(y+\alpha)>0$. Contradiction. So $B/\mathfrak{m}B$ is a radicial quadratic extension of $k$ and $B$ is regular.
If $\lambda_{\mathfrak{m}}(B)\geq 1$, then $a,b\in\mathfrak{m}$ and $B/\mathfrak{m}B\simeq k[T]/(T^{2})$. This implies that $B$ is local with maximal ideal $\mathfrak{n}$, not étale over $A$, and we have $k\simeq B/\mathfrak{n}$. The equality $B_{\mathfrak{n}}=B_{\mathfrak{m}}$ comes from the fact that for any $u\in B\setminus\mathfrak{n}$, the norm $N_{B/A}(u)\in A\setminus\mathfrak{m}$.
Finally when $\lambda_{\mathfrak{m}}(B)\geq 1$, we have $y\in\mathfrak{n}$ and the relation $y^{2}+ay+b=0$ implies that $B$ is regular if and only if $b\notin\mathfrak{m}^{2}$, equivalently, $\lambda_{\mathfrak{m}}(B)=1$. ∎
2.2. Algorithm for determining the multiplicity $\lambda_{p}(Y)$
We now discuss how to find a generator $y$ of $B$ reaching $\lambda_{\mathfrak{m}}(B)$. We can restrict to $\mathrm{char}(k)=2$ by Proposition 2.6. The lemma below is a generalization of [10], Lemma 3.3(1-2). It gives an algorithm to finding a generator $y$ of $B$ with $\lambda_{\mathfrak{m}}(B)=\lambda_{\mathfrak{m}}(y)$.
Lemma 2.10.
Let $A,\mathfrak{m},B$ be as in Definition 2.1 with $\mathrm{char}(A/\mathfrak{m})=2$. We omit $\mathfrak{m}$ from the subscripts of $\operatorname{ord}$ and $\lambda$.
  1. (1)
    If $2\operatorname{ord}(a)\leq\operatorname{ord}(b)$, then $\lambda(B)=\lambda(y)$.
  2. (2)
    Suppose $2\operatorname{ord}(a)>\operatorname{ord}(b)$. Let $\tilde{b}\in\mathrm{Gr}_{\mathfrak{m}}(A)$ be the initial form of $b$ (see Definition 2.1).
    1. (a)
      If $\tilde{b}$ is not a square in $\mathrm{Gr}_{\mathfrak{m}}(A)$, then $\lambda(B)=\lambda(y)$.
    2. (b)
      Suppose $\tilde{b}$ is a square in $\mathrm{Gr}_{\mathfrak{m}}(A)$: there exists $c\in A$ such that $b-c^{2}\in\mathfrak{m}^{\ell+1}$ where $\ell=\operatorname{ord}(b)$. Then
      $\lambda(y+c)>\lambda(y).$
Proof.
Up to multiplication by a unit of $A$, any other basis $\{1,z\}$ of $B$ over $A$ is given by $y=z+c$ for some $c\in A$. The equation of $z$ is
$z^{2}+(a+2c)z+(b+ac+c^{2})=0.$
Using the fact that $\operatorname{ord}$ is a valuation on $\mathrm{Frac}(A)$, the proof of (1) is similar to that of Lemma 3.3(1) in [10].
(2.a) If $\lambda(B)>\lambda(y)$, there exists $c\in A$ such that
$\operatorname{ord}(b+ac+c^{2})>\lambda(y)=\operatorname{ord}(b).$
As $2\operatorname{ord}(a)>\operatorname{ord}(b)$, this implies that $\operatorname{ord}(b)=2\operatorname{ord}(c)$ and $\operatorname{ord}(b+c^{2})>\operatorname{ord}(b)$. This is equivalent to the initial form of $b$ being a square in $\mathrm{Gr}_{\mathfrak{m}}(A)$.
(2.b) We have $\operatorname{ord}(b)=2\operatorname{ord}(c)$. Moreover
$2\operatorname{ord}(a+2c)>\operatorname{ord}(b),\quad\operatorname{ord}(b+ac+c^{2})=\operatorname{ord}(b-c^{2}+(a+2c)c)>\operatorname{ord}(b).$
Hence $\lambda(y+c)>\operatorname{ord}(b)=\lambda(y)$. ∎
  • Remark 2.11
    If $\mathrm{char}(A/\mathfrak{m})=2$, and $\Delta_{B/A}\neq 0$, then $\lambda_{\mathfrak{m}}(B)=\operatorname{ord}_{\mathfrak{m}}(\Delta_{B/A})$ if and only if for some generator $y$ of $B$ we have $2\operatorname{ord}_{\mathfrak{m}}(a)\leq\operatorname{ord}_{\mathfrak{m}}(b)$. This then implies that $\lambda_{\mathfrak{m}}(B)=\lambda_{\mathfrak{m}}(y)$.
  • Example 2.12
    Let $A=\mathbb{Z}[x]$. Let
    $B=\mathbb{Z}[x,y]/(y^{2}+x^{5}-1).$
    and let $\mathfrak{m}\subset A$ be a maximal ideal.
    1. (1)
      As $B\otimes_{\mathbb{Z}}\mathbb{Q}$ is regular, we have $\lambda_{\mathfrak{m}}(B)\leq 1$ if $\mathfrak{m}\cap\mathbb{Z}=\{0\}$.
    2. (2)
      Suppose now that $\mathfrak{m}\cap\mathbb{Z}=\ell\mathbb{Z}$ for some prime number $\ell$. There exists a monic $f(x)\in\mathbb{Z}[x]$ such that $f(x)$ is irreducible in $\mathbb{F}_{\ell}[x]$ and $\mathfrak{m}=(\ell,f(x))$.
      1. (a)
        If $\ell\neq 2,5$, then $\lambda_{\mathfrak{m}}(B)\leq 1$, with equality if and only if $f(x)\mid x^{5}-1$ in $\mathbb{F}_{\ell}[x]$.
      2. (b)
        If $\ell=5$, then $\lambda_{\mathfrak{m}}(B)=0$ if $f(x)$ is prime to $x-1$ in $\mathbb{F}_{5}[x]$. Let $\mathfrak{m}_{5}:=(5,x-1)$. Then $\lambda_{\mathfrak{m}_{5}}(B)=\operatorname{ord}_{\mathfrak{m}_{5}}(x^{5}-1)=2$.
      3. (c)
        Let $\ell=2$. Let $z=y-1\in B$. Then $z^{2}+2z+x^{5}=0$. Consider $\mathfrak{m}_{2}=(2,x)$. By Lemma 2.10(1), we have $\lambda_{\mathfrak{m}_{2}}(B)=2\operatorname{ord}_{\mathfrak{m}_{2}}(2)=2$.
      4. (d)
        Let $\ell=2$ and $\mathfrak{m}\neq\mathfrak{m}_{2}$, then $\lambda_{\mathfrak{m}}(B)=1$. Indeed, the curve $z^{2}+x^{5}$ over $\mathbb{F}_{2}$ is smooth away from the point $x=z=0$, hence $\operatorname{Spec}B\to\operatorname{Spec}\mathbb{Z}$ is smooth at $\mathfrak{m}$ and $B_{\mathfrak{m}}$ is regular. Thus $\lambda_{\mathfrak{m}}(B)=1$ (it is non-zero by Proposition 2.9(2)).
Proposition 2.13.
Let $(A,\mathfrak{m})\to(A^{\prime},\mathfrak{m}^{\prime})$ be an étale extension of regular local rings. Then
$\lambda_{\mathfrak{m}}(B)=\lambda_{\mathfrak{m}^{\prime}}(B\otimes_{A}A^{\prime}).$
Proof.
Let $k,k^{\prime}$ be the respective residue fields of $A,A^{\prime}$. As $\mathrm{Gr}_{\mathfrak{m}}(A)\to\mathrm{Gr}_{\mathfrak{m}^{\prime}}(A^{\prime})\simeq\mathrm{Gr}_{\mathfrak{m}}(A)\otimes_{k}k^{\prime}$ is injective, $\operatorname{ord}_{\mathfrak{m}}$ coincides with $\operatorname{ord}_{\mathfrak{m}^{\prime}}$ on $A$.
Let $y\in B$ such that $\lambda_{\mathfrak{m}}(y)=\lambda_{\mathfrak{m}}(B)$. Then $B^{\prime}=A^{\prime}[y]$. If $\mathrm{char}(k)\neq 2$ then $\lambda_{\mathfrak{m}^{\prime}}(B^{\prime})=\operatorname{ord}_{\mathfrak{m}^{\prime}}(\Delta_{B/A})=\operatorname{ord}_{\mathfrak{m}}(\Delta_{B/A})$. Suppose $\mathrm{char}(k)=2$. By Lemma 2.10, $\lambda_{\mathfrak{m}}(y)=\lambda_{\mathfrak{m}^{\prime}}(y)=\lambda_{\mathfrak{m}^{\prime}}(B^{\prime})$ (in the case 2.10(2) use the fact that if an element of $\mathrm{Gr}_{\mathfrak{m}}(A)$ is a square in $\mathrm{Gr}_{\mathfrak{m}}(A)\otimes_{k}k^{\prime}$, then it is already a square in $\mathrm{Gr}_{\mathfrak{m}}(A)$ because $k^{\prime}$ is separable over $k$.) ∎
2.3. Multiplicity in the global case
  • Definition 2.14
    Let $Z$ be an integral noetherian excellent regular scheme. Let $\psi:Y\to Z$ be a double cover (Definition 1). Then for any $p\in Y$ with image $q\in Z$, define
    $\lambda_{p}(Y)=\lambda_{q}(Y)=\lambda_{\mathfrak{m}_{q}}(\psi_{*}\mathcal{O}_{Y}\otimes_{\mathcal{O}_{Z}}\mathcal{O}_{Z,q}).$
    We define the discriminant ideal $\Delta_{Y/Z}\subset\mathcal{O}_{Z}$ as the norm $N_{\psi_{*}\mathcal{O}_{Y}/\mathcal{O}_{Z}}$ of the annihilator ideal of $\Omega^{1}_{Y/Z}$. It coincides with the usual definition when $Z=\operatorname{Spec}A$ is affine.
  • Remark 2.15
    The multiplicity $\lambda_{p}(Y)$ is the same for all points $p\in\psi^{-1}(q)$, so there is no ambiguity to denote it by $\lambda_{q}(Y)$. Moreover, when $\lambda_{p}(Y)\geq 1$, then $\psi^{-1}(q)=\{p\}$ and $k(q)\to k(p)$ is an isomorphism (Proposition 2.9(3)).
3. Normalized blowing-up
For any scheme $Y$ and closed subscheme $Y_{0}\subset Y$, we call the normalized blowing-up of $Y$ along $Y_{0}$ the normalization of the blowing-up of $Y$ along $Y_{0}$. In this section we study the normalized blowing-up of $Y$ at a singular point when $Y$ is a normal double cover $\psi:Y\to Z$ of an integral noetherian excellent regular surface $Z$.
3.1. Blowing-up regular schemes along a point
Let $A$ be a ring and let $\mathfrak{m}\subset A$ be a maximal ideal. We recall below the explicit description of the blowing-up
$\mathrm{bl}_{\mathfrak{m}}:Z^{\prime}\to Z=\operatorname{Spec}A$
of $Z$ along $\mathfrak{m}$. Denote by $\widetilde{A}$ the Rees algebra $\oplus_{n\geq 0}\mathfrak{m}^{n}$. For any $s\in\mathfrak{m}\setminus\{0\}$, denote by $\tilde{s}$ the element $s$ viewed as a homogeneous element of degree 111 in $\widetilde{A}$ and by $\widetilde{A}_{(\tilde{s})}$ the homogeneous localization of $\widetilde{A}$ with respect to the positive powers of $\tilde{s}$. Denote by $D_{+}(\tilde{s})\subseteq Z^{\prime}$ the principal open subset defined by $\tilde{s}$.
Proposition 3.1.
Let $A$ be an integral noetherian regular ring, let $\mathfrak{m}\subset A$ be a maximal ideal of $\dim A_{\mathfrak{m}}=2$. Let $s,t$ be a system of generators of $\mathfrak{m}$ and denote $k=A/\mathfrak{m}$.
  1. (1)
    We have
    $Z^{\prime}\simeq\operatorname{Proj}A[S,T]/(tS-sT),$
    $Z^{\prime}$ is a regular surface, covered by the principal open subsets $D_{+}(\tilde{s})$ and $D_{+}(\tilde{t})$. The homomorphism
    $A[T_{1}]/(sT_{1}-t)\to\widetilde{A}_{(\tilde{s})}=\mathcal{O}_{Z^{\prime}}(D_{+}(\tilde{s})),\quad T_{1}\mapsto\tilde{t}/\tilde{s}$
    of $A$-algebras is an isomorphism, it takes $\mathfrak{m}^{r}\subseteq A$ onto $s^{r}\widetilde{A}_{(\tilde{s})}$.
  2. (2)
    Let $E=V(\mathfrak{m}\mathcal{O}_{Z^{\prime}})\subset Z^{\prime}$ be the exceptional divisor. Then
    $\cap D_{+}(\tilde{s})=V(s\widetilde{A}_{(\tilde{s})})\simeq\operatorname{Spec}k[\bar{t}_{1}]$
    where $\bar{t}_{1}$ is the images of $t_{1}:=\tilde{t}/\tilde{s}$ modulo $s$. Any maximal ideal of $\widetilde{A}_{(\tilde{s})}$ containing $s$ is generated by $s$ and one other element.
  3. (3)
    Let $\nu_{E}$ be the normalized discrete valuation on $\mathrm{Frac}(A)$ associated to the generic point of $E$. Then $\nu_{E}(\alpha)=\operatorname{ord}_{\mathfrak{m}}(\alpha)$ for all $\alpha\in\mathrm{Frac}(A)$.
  4. (4)
    If $A$ is factorial then so is $\widetilde{A}_{(\tilde{s})}$.
  5. (5)
    Let $f\in A$ with $\operatorname{ord}_{\mathfrak{m}}(f)=n\geq 1$.
    1. (a)
      Let $V(f)^{\prime}$ be the strict transform of $V(f)$ in $Z^{\prime}$. Then we have $V(f)^{\prime}\cap D_{+}(\tilde{s})=V(f/s^{n})$.
    2. (b)
      The image of $f/s^{n}\in\widetilde{A}_{(\tilde{s})}$ in $\widetilde{A}_{(\tilde{s})}/(s)\simeq k[\bar{t}_{1}]$ has degree at most equal to $n$.
Proof.
(1), (2), (5.a) are standard.
(3) Let $\xi_{E}$ be the generic point of $E$. By (2), $\xi_{E}\in D_{+}(\tilde{s})$ and $\mathcal{O}_{Z^{\prime},\xi_{E}}(-E)$ is generated by $s$. Therefore $\nu_{E}(s)=1$. If $\alpha\in\mathfrak{m}^{\ell}\setminus\mathfrak{m}^{\ell+1}$, then $\alpha/s^{\ell}\in\widetilde{A}_{(\tilde{s})}$, hence $\nu_{E}(\alpha)\geq\ell=\operatorname{ord}_{\mathfrak{m}}(\alpha)$.
Conversely, suppose $\nu_{E}(\alpha)=\ell$. We have
$\mathrm{div}(\alpha s^{-\ell})=\mathrm{div}(\alpha)-\ell E\geq 0$
as Weil divisors on $D_{+}(\tilde{s})$. Thus $\alpha\in s^{\ell}\widetilde{A}_{(\tilde{s})}$. So there exists $N\geq\ell$ such that $\alpha s^{N-\ell}\in\mathfrak{m}^{N}$. Because the class of $s$ in $\mathfrak{m}/\mathfrak{m}^{2}\subset\mathrm{Gr}_{\mathfrak{m}}(A)$ is a regular element, $\alpha\in\mathfrak{m}^{\ell}$, hence $\operatorname{ord}_{\mathfrak{m}}(\alpha)\geq\nu_{E}(\alpha)$.
(4) Notice that $s\in\widetilde{A}_{(\tilde{s})}$ is a prime element by (2), and $(\widetilde{A}_{(\tilde{s})})_{s}=A_{s}$ is factorial. Hence $\widetilde{A}_{(\tilde{s})}$ is factorial by Nagata’s criterion.
(5.b) Write
$f=\sum_{0\leq i\leq n}\alpha_{i}t^{i}s^{n-i}+\epsilon_{n+1}$
with $\alpha_{i}\in A$ and $\epsilon_{n+1}\in\mathfrak{m}^{n+1}$. This implies that
$f/s^{n}=\sum_{0\in i\leq n}\alpha_{i}t_{1}^{i}+sc,\quad c\in\widetilde{A}_{(\tilde{s})}.$
Its class mod $s$ has degree $\leq n$. ∎
3.2. Normalization after blowing-up
To determine the integral closure of an integral domain we will use Serre’s (S2) + (R1) criterion ([9], Theorem 8.2.23).
Lemma 3.2.
Let $(R,\nu)$ be a discrete valuation ring with residue field $\kappa$, let
$D=R[y]/(y^{2}+\alpha y+\beta)$
be an $R$-algebra with $\Delta_{D/R}:=\alpha^{2}-4\beta\neq 0$.
  1. (1)
    If $\nu(\alpha^{2}-4\beta)=0$, then $D$ is an étale $R$-algebra of rank 222;
  2. (2)
    If $\nu(\alpha^{2}-4\beta)=1$, then $D$ is a discrete valuation ring totally ramified over $R$ with ramification index 222;
  3. (3)
    The same is true if $\mathrm{char}(\kappa)=2$, $\nu(\alpha)>0$, and $\nu(\beta)=1$;
  4. (4)
    If $\mathrm{char}(\kappa)=2$, $\nu(\alpha)>0$, and $\beta$ is not a square in $\kappa$, then $D$ is a discrete valuation ring with ramification index 111 over $R$ and inseparable quadratic residue extension.
Proof.
Straightforward. ∎
Proposition 3.3.
Keep the notation of Proposition 3.1 with $A$ excellent. Let $B$ be an integrally closed domain, free of rank 222 over $A$. Fix a generator $y$ such that $\lambda_{\mathfrak{m}}(y)=\lambda_{\mathfrak{m}}(B)$ and set $r=[\lambda_{\mathfrak{m}}(B)/2]$. Let $s\in\mathfrak{m}\setminus\mathfrak{m}^{2}$ and let $B_{1}:=\widetilde{A}_{(\tilde{s})}[y_{1}]\subset\mathrm{Frac}(B)$. Put
$y_{1}=y/s^{r},\ a_{1}=a/s^{r},\ b_{1}=b/s^{2r}\in A_{s}.$
Then
  1. (1)
    $\Delta_{B_{1}/\widetilde{A}_{(\tilde{s})}}=s^{-2r}\Delta_{B/A}\widetilde{A}_{(\tilde{s})}$;
  2. (2)
    $a_{1},b_{1}\in\widetilde{A}_{(\tilde{s})}$,
    $y_{1}^{2}+a_{1}y_{1}+b_{1}=0$ (3.1)
    and $B_{1}$ is the integral closure of $\widetilde{A}_{(\tilde{s})}$ in $\mathrm{Frac}(B)$.
Proof.
The hypothesis $A$ excellent insures that $\lambda_{\mathfrak{m}}(B)$ is finite (Proposition 2.6). Part (2) is immediate.
(1) We have $a_{1},b_{1}\in\widetilde{A}_{(\tilde{s})}$ by Proposition 3.1(1). As $B_{1}$ contains $\widetilde{A}_{(\tilde{s})}$ and is finite over $\widetilde{A}_{(\tilde{s})}$, it is enough to show that $B_{1}$ is integrally closed. As $\widetilde{A}_{(\tilde{s})}$ is regular and
$B_{1}\simeq\widetilde{A}_{(\tilde{s})}[T]/(T^{2}+a_{1}T+b_{1}),$
$B_{1}$ is a local complete intersection, hence satisfies (Sn) for all $n\geq 1$ ([9], Corollary 8.2.18 and Example 8.2.20). By Serre’s criterion it is enough to show that $B_{1}$ is integrally closed at points of codimension 111. Moreover, as $\operatorname{Spec}B_{1}\to\operatorname{Spec}B$ is an isomorphism away from $V(s)$, it is enough to check the property for codimension 111 points belonging to $V(s)$, so lying over the generic point $\xi_{E}$ of $E$ (exceptional divisor of $Z^{\prime}\to Z$).
Let $R=\mathcal{O}_{Z^{\prime},\xi_{E}}$. Then
$B_{1}\otimes_{\widetilde{A}_{(\tilde{s})}}R\simeq R[y_{1}]/(y_{1}^{2}+a_{1}y_{1}+b_{1}).$
We will apply Lemma 3.2. We have $\min\{2\nu_{E}(a_{1}),\nu_{E}(b_{1})\}\leq 1$. So we only have to check that when $\mathrm{char}(k(\xi_{E}))=2$, $\nu_{E}(a_{1})>0$ and $\nu_{E}(b_{1})=0$, then $b_{1}$ is not a square in $k(\xi_{E})$. Under these hypothesis, we have $2\operatorname{ord}_{\mathfrak{m}}(a)=2\nu_{E}(a)>2r=\nu_{E}(b)=\operatorname{ord}_{\mathfrak{m}}(b)$ (Proposition 3.1(3)).
Suppose that $b_{1}$ is a square in $k(\xi_{E})=\mathrm{Frac}(\widetilde{A}_{(\tilde{s})}/(s))$. As $\widetilde{A}_{(\tilde{s})}/(s)$ is integrally closed (Proposition 3.1(2)), $b_{1}$ is a square in $\widetilde{A}_{(\tilde{s})}/(s)$: there exist $c_{1},c_{2}\in\widetilde{A}_{(\tilde{s})}$ such that
$b_{1}=c_{1}^{2}+sc_{2}.$
There exist $N\geq r$ big enough, $\alpha_{N}\in\mathfrak{m}^{N}$ and $\beta_{2N}\in\mathfrak{m}^{2N}$ such that $c_{1}=\alpha_{N}/s^{N}$, $c_{2}=\beta_{2N}/s^{2N}$. Then
$s^{2N-2r}b=\alpha_{N}^{2}+s\beta_{2N}.$
So the initial form of $s^{2N-2r}b\in\mathfrak{m}^{2N}$ is a square in $\mathrm{Gr}_{\mathfrak{m}}(A)$, thus the same is true for that of $b$. This contradicts Lemma 2.10(2.b) and (1) is proved. ∎
3.3. Lipman’s resolution of singularities
Theorem 3.4.
Let $Z$ be an integral noetherian excellent regular surface, let $\psi:Y\to Z$ be a normal double cover. Let $p_{0}\in Y$ be a singular point, let $q_{0}=\psi(p_{0})$.
  1. (1)
    Let $Z^{\prime}$ (resp. $Y^{\prime}$) be the blowing-up of $Z$ (resp. of $Y$) along the (reduced) closed point $q_{0}$ (resp. $p_{0}$) and let $\tilde{Y}\to Y^{\prime}$ be the normalization. Then we have a canonical commutative diagram
[…]
    and $\phi:\tilde{Y}\to Z^{\prime}$ is a double cover with discriminant ideal
    $\Delta_{\tilde{Y}/Z^{\prime}}=\mathcal{O}_{Z^{\prime}}(2rE)\otimes_{\mathcal{O}_{Z}}\Delta_{Y/Z},$
    where $r=[\lambda_{q_{0}}(Y)/2]$ and $E$ is the exceptional divisor of $Z^{\prime}\to Z$.
  2. (2)
    If
    $Y_{n}\to Y_{n-1}\to...\to Y_{1}\to Y_{0}=Y$
    is a sequence of normalized blowing-ups of (reduced) singular points $p_{i}\in Y_{i}$, then we have a commutative diagram
[…]
    where $Z_{i+1}\to Z_{i}$ is the blowing-up of the reduced closed point $q_{i}\in Z_{i}$, image of $p_{i}$, and each $\psi_{i}:Y_{i}\to Z_{i}$ is a normal double cover.
  3. (3)
    Suppose moreover that $Y\to Z$ is generically étale. Then the Galois group $G:=\mathrm{Gal}(K(Y)/K(Z))$ acts on each $Y_{i}$ and $Y_{i}/G=Z_{i}$.
Proof.
(1) First we prove that $Y^{\prime}\to Z$ factorizes through a finite morphism $Y^{\prime}\to Z^{\prime}$. We can suppose that $B$ is free over $A$. Let $\mathfrak{m}\subset A$ be the maximal ideal corresponding to $q_{0}$. Let $\lambda_{\mathfrak{m}}(y)=\lambda_{\mathfrak{m}}(B)$. Then $a\in\mathfrak{m}$ and $b\in\mathfrak{m}^{2}$. Let $\mathfrak{n}\subset B$ be the maximal ideal corresponding to $p_{0}$. It is easy to that
$\mathfrak{n}^{n}=\mathfrak{m}^{n}\oplus\mathfrak{m}^{n-1}y=\mathfrak{m}^{n-1}\mathfrak{n}$
for all $n\geq 1$. Therefore
$\oplus_{n\geq 0}\mathfrak{n}^{n}=Ay+(\oplus_{n\geq 0}\mathfrak{m}^{n})+(\oplus_{n\geq 0}\mathfrak{m}^{n})\tilde{y}$
where $\tilde{y}$ is $y$ viewed as homogeneous element of degree 111. Hence $\oplus_{n\geq 0}\mathfrak{n}^{n}$ is finite over $\oplus_{n\geq 0}\mathfrak{m}^{n}$ (see also [20], Proposition 5.2.5), and $Y\to Z$ induces a finite morphism $Y^{\prime}\to Z^{\prime}$. Thus $\tilde{Y}$ coincides with the normalization of $Z^{\prime}$ in $K(Y)$. The local structure of $\tilde{Y}$ is described by Proposition 3.3, and we see that $\tilde{Y}\to Z^{\prime}$ is a normal double cover (this also follows from Example 1), and the discriminant ideal is $\mathcal{O}_{Z^{\prime}}(2rE)\otimes_{\mathcal{O}_{Z}}\Delta_{Y/Z}$.
(2) and (3) follow from (1). ∎
  • Remark 3.5
    Suppose $Z=\operatorname{Spec}A$ is affine and $Y=\operatorname{Spec}B$ is free over $A$. Consider the graded $A$-algebra $\tilde{A}[W]$ with $\deg W=r$. We see using Proposition 3.3 that $\tilde{Y}$ is the closed subscheme of $\operatorname{Proj}(\tilde{A}[W])$ defined by the equation
    $W^{2}+\tilde{a}W+\tilde{b}=0,$
    where $\tilde{a},\tilde{b}$ are respectively $a$ and $b$ considered as homogeneous elements in $\tilde{A}$ of respective degrees $r,2r$.
3.4. The exceptional locus of Lipman’s resolution of singularities
Keep the assumption and notation of Theorem 3.4. Suppose that Lipman’s sequence for the singular point $p_{0}\in Y_{0}$ terminates at $Y_{N}$. The morphism $Y_{N}\to Y_{0}$ is an isomorphism away from $p_{0}$. In this subsection we study the scheme structure of the pre-image of $p_{0}$ by $Y_{N}\to Y_{0}$ (the exceptional locus of $Y_{N}\to Y_{0}$).
We fix a prime divisor $E_{0}\ni q_{0}$ of $Z_{0}$, regular at $q_{0}$. Typically, for our applications in mind, $Z$ is a relative curve over a discrete valuation ring $R$, and $E_{0}$ is the closed fiber of $Z$. Let
$\mathcal{E}:=Y_{N}\times_{Y_{0}}\psi_{0}^{*}E_{0}\subset Y_{N}.$
This is the union of the strict transform of $\psi^{*}E_{0}$ in $Y_{N}$ and the exceptional locus of $Y_{N}\to Y_{0}$. In the above example, $\mathcal{E}$ is the closed fiber of $Y_{N}$. We would like to know the structure of each irreducible component of $\mathcal{E}$, their multiplicities in $\mathcal{E}$, and how they intersect each other.
The double cover $Y_{0}\to Z_{0}$ induces a finite flat morphism $\mathcal{E}\to Z_{N}\times_{Z_{0}}E_{0}$ of degree 222. We start by describing $Z_{N}\times_{Z_{0}}E_{0}$. For all $1\leq i\leq N$, denote by $E_{i}\subset Z_{i}$ the exceptional divisor of $Z_{i}\to Z_{i-1}$ and by $E^{\prime}_{i}$ (including $i=0$) its strict transform in $Z_{N}$.
Lemma 3.6.
Write
$Z_{N}\times_{Z}E_{0}=\sum_{0\leq i\leq N}m_{i}E^{\prime}_{i}$
as Cartier divisors of $Z_{N}$.
  1. (1)
    For $1\leq i\leq N$, $E_{i}=\mathbb{P}^{1}_{k(q_{i-1})}$. For all $0\leq i\leq N$, the canonical morphism $E^{\prime}_{i}\to E_{i}$ is an isomorphism. Denote by $q^{\prime}_{i}\in E^{\prime}_{i}$ the pre-image of $q_{i}$.
  2. (2)
    The union $\cup_{0\leq i\leq N}E^{\prime}_{i}\subset Z_{N}$ is a tree of irreducible components with transverse intersections.
  3. (3)
    The multiplicities $m_{i}$ can be determined inductively as follows: $m_{0}=m_{1}=1$. Let $i\geq 1$. If $q_{i}$ is the intersection point of $E_{i}$ with the strict transform of $E_{i-1}$ in $Z_{i}$, then $m_{i+1}=m_{i}+m_{i-1}$. Otherwise, $m_{i+1}=m_{i}$.
Proof.
(1) This is because $E^{\prime}_{i}\to E_{i}$ is finite birational and $E_{i}$ is normal.
(2)-(3) Apply inductively Proposition 3.1 to the $Z_{i}$’s. One can take for $s$ a generator of $\mathcal{O}_{Z}(-E_{0})_{q_{0}}$. ∎
Proposition 3.7.
Keep the notation of Lemma 3.6. For $1\leq i\leq N$, set
$\Theta^{\prime}_{i}=\psi_{N}^{-1}(E^{\prime}_{i})$
with the reduced structure.
  1. (1)
    The scheme $\Theta^{\prime}_{i}$ is irreducible if and only if $\psi_{i}^{-1}(E_{i})$ is irreducible. When it is reducible, it has two irreducible components, each of them is isomorphic to $E^{\prime}_{i}$ by $\psi_{N}$, and has multiplicity $m_{i}$ in $\mathcal{E}$.
  2. (2)
    Suppose that $\lambda_{q_{i-1}}(Y_{i-1})$ is odd, then $\Theta^{\prime}_{i}\to E^{\prime}_{i}$ is an isomorphism, and the multiplicity of $\Theta^{\prime}_{i}$ in $\mathcal{E}$ is $2m_{i}$.
  3. (3)
    Suppose that $\lambda_{q_{i-1}}(Y_{i-1})$ is even and $\Theta^{\prime}_{i}$ is irreducible. Then the latter has multiplicity $m_{i}$ in $\mathcal{E}$. Moreover, if $\mathrm{char}(k(q_{0}))\neq 2$, then $\Theta^{\prime}_{i}\to E^{\prime}_{i}$ is generically étale.
Proof.
Note that any finite birational morphism onto $E^{\prime}_{i}$ is an isomorphism because the latter is normal. So it is enough to study $\Theta^{\prime}_{i}$ at its generic points. Note that the multiplicity in $\mathcal{E}$ of an irreducible component of $\Theta^{\prime}_{i}$ is equal to $m_{i}$ multiplied by the ramification index over $E^{\prime}_{i}$.
As $Y_{N}\to Y_{i}$ is an isomorphism over an open neighborhood of the generic point of $E_{i}$, it is enough to show the statement for $\Theta_{i}:=\psi_{i}^{-1}(E_{i})$. Moreover without loss of generalities we can assume that $i=1$. Then all the statements of the proposition follow from direct observations on an affine equation (3.1) of $Y_{1}$
  • Remark 3.8
    Keep the notation of Propositions 3.7 and 3.3.
    1. (1)
      The scheme $\psi_{1}^{-1}(E_{1})$ is irreducible if and only if the polynomial
      $T^{2}+\bar{a}_{1}T+\bar{b}_{1}\in k[\bar{t}_{1}][T]$
      (where $\bar{*}$ means the class modulo $s$) is irreducible. The same statement applies to $\psi_{i}^{-1}(E_{i})$ if a similar equation of $Y_{i}$ is known.
    2. (2)
      Suppose $\mathrm{char}(k(q_{0}))=2$ and $\Theta_{1}$ is irreducible. The $\Theta^{\prime}_{1}\to E^{\prime}_{1}$ is generically étale if and only if $2\operatorname{ord}_{\mathfrak{m}}(a)=\lambda_{q_{0}}(Y)$.
    3. (3)
      The structure of $\Theta^{\prime}_{1}$, in the case (3) of Proposition 3.7 is more complicate to describe. We have to determine inductively the strict transform of $\Theta_{1}$ after each blowing-up.
Proposition 3.9.
Keep the notation of Lemma 3.6 and Proposition 3.7. Let $0\leq i\neq j\leq N$ be such that $E^{\prime}_{i}\cap E^{\prime}_{j}=\{q\}\neq\emptyset$.
  1. (1)
    If $\Theta^{\prime}_{i}$ has two irreducible components $\Theta^{\prime}_{i,1},\Theta^{\prime}_{i,2}$ and similarly for $\Theta^{\prime}_{j}$, then up to renumbering, $\Theta^{\prime}_{i,1}$ and $\Theta^{\prime}_{j,1}$ (resp. $\Theta^{\prime}_{i,2}$ and $\Theta^{\prime}_{j,2}$) intersect transversely at a single $k(q)$-rational point, and $\Theta^{\prime}_{i,1}\cap\Theta^{\prime}_{j,2}=\Theta^{\prime}_{i,2}\cap\Theta^{\prime}_{j,1}=\emptyset$.
  2. (2)
    If $\Theta^{\prime}_{i}$ has two irreducible components and $\Theta^{\prime}_{j}$ is irreducible, then $\Theta^{\prime}_{j}$ intersects transversely each of these components at a single $k(q)$-rational point. These intersection points can be the same.
  3. (3)
    Suppose that $\Theta^{\prime}_{i}$ and $\Theta^{\prime}_{j}$ are both irreducible. Denote by $e_{i},e_{j}$ the respective ramification indexes of $\psi_{N}$ at the generic points of $\Theta^{\prime}_{i},\Theta^{\prime}_{j}$. Then we have
    $e_{i}[k(\Theta^{\prime}_{i}):k(E^{\prime}_{i})]=e_{j}[k(\Theta^{\prime}_{j}):k(E^{\prime}_{j})]=2$
    and
    $\sum_{p\in\psi_{N}^{-1}(q)}i_{p}(\Theta^{\prime}_{i},\Theta^{\prime}_{j})e_{i}e_{j}[k(p):k(q)]=2$
    where $i_{p}(\Theta^{\prime}_{i},\Theta^{\prime}_{j})$ denotes the intersection multiplicity of $\Theta^{\prime}_{i}$ and $\Theta^{\prime}_{j}$ at $p$.
Proof.
Let $\{\Theta^{\prime}_{i,\alpha}\}_{\alpha},\{\Theta^{\prime}_{j,\beta}\}_{\beta}$ be the respective irreducible components of $\Theta^{\prime}_{i}$ and $\Theta^{\prime}_{j}$. Let $e_{i,\alpha}$ be the ramification index of $\psi_{N}$ at the generic point of $\Theta^{\prime}_{i,\alpha}$, similarly for $e_{j,\beta}$. We have
$\sum_{\alpha}e_{i,\alpha}[k(\Theta^{\prime}_{i,\alpha}):k(E^{\prime}_{i})]=[K(Y):K(Z)]=2.$
Similarly for $\Theta^{\prime}_{j}$. By general intersection theory, we have
$\sum_{p\in\psi_{N}^{-1}(q)}i_{p}(\psi_{N}^{*}E^{\prime}_{i},\psi_{N}^{*}E^{\prime}_{j})[k(p):k(q)]=[K(Y):K(Z)]i_{q}(E^{\prime}_{i},E^{\prime}_{j})=2,$
and by definition $\psi_{N}^{*}E^{\prime}_{i}=\sum_{\alpha}e_{i,\alpha}\Theta^{\prime}_{i,\alpha}$ (similarly for $\psi_{N}^{*}E^{\prime}_{j}$). Therefore
$\sum_{p\in\psi_{N}^{-1}(q)}\sum_{\alpha,\beta}i_{p}(\Theta^{\prime}_{i,\alpha},\Theta^{\prime}_{j,\beta})e_{i,\alpha}e_{j,\beta}[k(p):k(q)]=2.$
All coefficients in the above equality are positive except the intersection multiplicities $i_{p}$ which can be zero. This leaves very few possibilities for the combinations of the various coefficients.
We also notice that through any point of $\psi_{N}^{-1}(q)\cap\Theta^{\prime}_{i}$ passes a point of $\psi_{N}^{-1}(q)\cap\Theta^{\prime}_{j}$, by the going down property of $Y_{N}\to Z_{N}$ (see [19], tag 00HU). The proposition is then proved by straightforward case-by-case analysis. ∎
3.5. A complete example of resolution of singularities
  • Example 3.10
    Let $Z=\mathbb{P}^{1}_{\mathbb{Z}_{2}}$ be the projective line over the 222-adic integers. Fix two integers $n,m$ such that at least one of them is non-zero. Let $Y$ be the projective scheme over $\mathbb{Z}_{2}$ associated to the affine equation
    $y^{2}=(s^{2}-2)^{3}+16ms^{2}+n^{2}.$
    Replacing $y$ with $y+s^{3}+n$, we get
    $y^{2}+2(s^{3}+n)y+2(3s^{4}+ns^{3}+2m_{1}s^{2}+4)=0$ (3.2)
    where $m_{1}:=-(4m+3)$. This is a normal Weierstrass model of a genus 2 curve ([10], Lemma 2.3.1(c)). The closed fiber of $Y$ over $\operatorname{Spec}\mathbb{Z}_{2}$ is a double (multiplicity 222) projective line over $\mathbb{F}_{2}$. Denote it by $2\Theta_{0}$. The aim of this example is to determine the minimal resolution of singularities of $Y$ when $8\nmid n$. Denote by $\boxed{t=2}$ the uniformizing element of $\mathbb{Z}_{2}$.
    For each singular point of $Y$, the notations $Y_{i}$ and $Z_{i}$ refer to the blowing-ups described by Theorem 3.4. We use the same notation for different singular points for simplicity. When we describe the equations of $Y_{i}$, $D_{+}(\tilde{f})$ always means open subset in $Z_{i+1}$ (for instance, we frequently consider $D_{+}(\tilde{s})$ as open subset of $Z_{i}$ for $i>1$).
[…]
    $z^{2}+2(1+nx^{3})z+2x^{2}(3+nx+2m_{1}x^{2}+4x^{4})=0.$ (3.3)
    with $\lambda_{q_{\infty}}(Y)=2=\lambda_{q_{\infty}}(z)$.
    (I.1) First blowing-up. We blow-up $Y_{1}\to Y$ along the point lying over $q_{\infty}$. Over $D_{+}(\tilde{x})$, we have $t=xt_{1}$ and
    $z_{1}^{2}+t_{1}(1+nx^{3})z_{1}+t_{1}x(3+nx+2m_{1}x^{2}+4x^{4})=0.$ (3.4)
    There is a unique singular point $p_{\infty,1}=(t_{1}=x=z_{1}=0)$ with multiplicity $\lambda_{p_{\infty,1}}(Y_{1})=2$.
    Over $D_{+}(\tilde{t})$, the equation is ($x_{1}=x/t$)
    $w_{1}^{2}+(1+nt^{3}x_{1}^{3})w_{1}+tx_{1}^{2}(3+ntx_{1}+2m_{1}t^{2}x_{1}^{2}+4t^{4}x_{1}^{4})=0$
    which defines two regular points at $x_{1}=0$ (poles of $t_{1}$).
    The exceptional locus $\Theta_{1}$ (with the reduced structure) of $Y_{1}\to Y$ is union of two projective lines over $\mathbb{F}_{2}$, of multiplicity 111, intersecting at $p_{\infty,1}$. The strict transform of $\Theta_{0}$ intersects $\Theta_{1}$ at $t_{1}=0$ which is $p_{\infty,1}$.
    (I.2) Second blowing-up. We blow-up $Y_{2}\to Y_{1}$ along $p_{\infty,1}$. We use Equation (3.4). Over $D_{+}(\tilde{x})\subset Z_{1}$ we have (letting $t_{1}=xt_{2}$)
    $z_{2}^{2}+t_{2}(1+nx^{3})z_{2}+t_{2}(3+nx+2m_{1}x^{2}+4x^{4})=0.$
    Over $D_{+}(\tilde{t}_{1})$, we have $x=t_{1}x_{1}$ and
    $w_{2}^{2}+(1+nt_{1}^{3}x_{1}^{3})w_{2}+x_{1}(3+nt_{1}x_{1}+2m_{1}t_{1}^{2}x_{1}^{2}+4t_{1}^{4}x_{1}^{4})=0.$
    We see that the exceptional locus is a smooth conic of multiplicity 222 (we have $t=x^{2}t_{2}$). Its intersection with the strict transform of $\Theta_{0}$ is the point $t_{2}=z_{2}=0$. The intersection with the strict transform of $\Theta_{1}$ are the points $x_{1}=0,w_{2}=0,1$.
    The exceptional locus of $Y_{1}\to Y$ and $Y_{2}\to Y_{1}$ as well as the strict transforms of $\Theta_{0}$ are as shown in the figure below. In $Y_{2}$, the components are isomorphic to $\mathbb{P}^{1}_{\mathbb{F}_{2}}$, the intersections are transverse at rational points.
    The numbers indicate the multiplicities of the components.
    (II) Now come back to Equation (3.2). The singularities depend on the parity of $n$. Indeed, when $n$ is even, there is a unique singular point $p_{0}$, lying above $q_{0}=(s=2=0)$ with $\lambda_{q_{0}}(Y)=3$. When $n$ is odd, there are two singular points: $p_{0}$ with $\lambda_{q_{0}}(Y)=2$ and $p_{*}$ lying over $q_{*}=(s+1=2=0)$, with $\lambda_{q_{*}}(Y)=2$.
    We start with the simpler case $\boxed{n\text{ is odd}}$.
(II.1) First blowing-up. We blow-up $Y_{1}\to Y$ along $p_{0}$. We use Equation (3.2). The equation of $Y_{1}$ over $D_{+}(\tilde{s})$ is ($2=t=st_{1}$):
$y_{1}^{2}+t_{1}(s^{3}+n)y_{1}+st_{1}(3s^{2}+ns+2m_{1}st_{1}+t_{1}^{2})=0.$ (3.5)
There is a unique singular point $p_{1}$, lying over $q_{1}=(s=t_{1}=0)$, with $\lambda_{q_{1}}(Y_{1})=2$. Over $D_{+}(\tilde{t})$, we have $s=s_{1}t$ and
$w_{1}^{2}+(s_{1}^{3}t^{3}+n)w_{1}+t(3s_{1}^{4}t^{2}+ns_{1}^{3}t+2m_{1}s_{1}^{2}+1)=0$
which is regular at $s_{1}=0$ (two points, poles of $t_{1}$).
The exceptional locus $\Theta_{1}$ of $Y_{1}\to Y$ above $p_{0}$ is a reduced degenerate conic as $n$ is odd. It meets the strict transform of $\Theta_{0}$ at $p_{1}$.
(II.2) Second blowing-up. We blow-up $Y_{2}\to Y_{1}$ along $p_{1}$. We use Equation (3.5). Over $D_{+}(\tilde{s})\subset Z_{2}$, we have $t_{1}=st_{2}$ and
$y_{2}^{2}+t_{2}(s^{3}+n)y_{2}+t_{2}s(n+3s+m_{1}st_{2}+st_{2}^{2})=0.$ (3.6)
Over $D_{+}(\tilde{t}_{1})$, the equation of $Y_{2}$ is ($s=s_{2}t_{1}$)
$w_{2}^{2}+(n+s_{2}^{3}t_{1}^{3})w_{2}+s_{2}t_{1}(3s_{2}^{2}t_{1}+ns_{2}+m_{1}s_{2}t_{1}+t_{1})=0,$
which is regular at $s_{2}=0$.
The exceptional locus $\Theta_{2}$ is union of two projective lines over $\mathbb{F}_{2}$, of multiplicity 222 and intersecting at $p_{2}=(t_{2}=y_{2}=s=0)$ with $\lambda_{p_{2}}(Y_{2})=2$. They have multiplicity 222 because $t=s^{2}t_{2}$. The intersection with the strict transform of $\Theta_{0}$ is $p_{2}$, and with that of $\Theta_{1}$ is $s_{2}=0$ (two points).
(II.3) Third blowing-up. We blow-up $Y_{2}$ along the unique singular point $p_{2}$ (lying over $p_{0}$). Over $D_{+}(\tilde{s})\subset Z_{3}$, we have ($t_{2}=st_{3}$)
$y_{3}^{2}+t_{3}(s^{3}+n)y_{3}+t_{3}(n+3s+m_{1}s^{3}t_{3}+s^{3}t_{3}^{2})=0.$
Over $D_{+}(\tilde{t}_{2})$, we have ($s=s_{3}t_{2}$)
$w_{3}^{2}+(s_{3}^{3}t_{2}^{3}+n)w_{3}+s_{3}(n+3s_{3}t_{2}+m_{1}s_{3}t_{2}^{2}+s_{3}t_{2}^{3})=0.$
The exceptional locus $\Theta_{3}$ is a smooth conic, intersecting the strict transform of $\Theta_{2}$ at the two rationals points $s_{3}=0$. It intersects the strict transform of $\Theta_{0}$ at the rational point $t_{3}=0$. As $t=s^{3}t_{3}$, $\Theta_{3}$ has multiplicity 333. We finished solving the singularity $p_{0}$. We represent the three blowing-ups above in the picture below.
In the last step all components are isomorphic to $\mathbb{P}^{1}_{\mathbb{F}_{2}}$ and the intersection points are rational and transverse.
(III) The resolution of $p_{*}$ is obtained by one single blowing-up. The exceptional locus consists in a copy of $\mathbb{P}^{1}_{\mathbb{F}_{2}}$ of multiplicity 111, meeting transversely the strict transform of $\Theta_{0}$ at a rational point.
The complete resolution of the projective Weierstrass model defined by Equation (3.2) is of type $[{\tt IV}^{*}-{\tt I}^{*}_{1}-\alpha]$ in [16], page 175.
(IV) Suppose from now on that $\boxed{n=2n_{1}\text{\ is even}}$. We already solved $p_{\infty}$ in Part (I). It remains to solve the singularity at $p_{0}$. We have $\lambda_{p_{0}}(Y)=3$ given by Equation (3.2).
(IV.1) First normalized blowing-up. We use Equation (3.2). Let $Y_{1}\to Y$ be the normalized blowing-up along $p_{0}$. The equation over $D_{+}(\tilde{s})$ is
$y^{2}_{1}+st_{1}(s^{2}+n_{1}t_{1})y_{1}+t_{1}(3s^{3}+n_{1}s^{3}t_{1}+m_{1}s^{2}t_{1}+st_{1}^{2})=0.$ (3.7)
The point at infinity, $s_{1}=0$ in $D_{+}(\tilde{t})$, is regular in $Y_{1}$. The exceptional locus $\Theta_{1}$ is $2\mathbb{P}^{1}_{\mathbb{F}_{2}}$. It intersects the strict transform of $\Theta_{0}$ at $p_{1}=(s=t_{1}=y_{1}=0)$, with $\lambda_{p_{1}}(Y_{1})=4$.
(IV.2) Normalized blowing-up along $p_{1}$. We use Equation (3.7). Over $D_{+}(\tilde{s})\subset Z_{1}$, we have $t_{1}=st_{2}$ and
$y_{2}^{2}+st_{2}(s+n_{1}t_{2})y_{2}+t_{2}(3+n_{1}st_{2}+m_{1}t_{2}+t_{2}^{2})=0.$ (3.8)
The point at infinity $s_{2}=0$ ($s=t_{1}s_{2}$), is regular in $Y_{2}$
The exceptional locus $\Theta_{2}$ is a rational curve over $\mathbb{F}_{2}$ of multiplicity 222, with a cusp $p_{2}$ at $t_{2}=1$. The intersection of $\Theta_{2}$ with the strict transform of $\Theta_{0}$ is $t_{2}=0$, and with that of $\Theta_{1}$ at $s_{2}=0$. The strict transforms of $\Theta_{0}$ and of $\Theta_{1}$ do not intersect each other.
The only possible singular point of $Y_{2}$ lying over $p_{0}$ is $p_{2}$. Let us compute $\lambda_{p_{2}}(Y_{2})$. We make the change of variables $t_{2}=x+1$ and $y_{2}=z_{2}+1$. We have
$z_{2}^{2}+n_{1}s(1+x)^{2}z_{2}+x^{3}-ms^{4}(1+x)^{4}=0.$ (3.9)
The strict transform of $\Theta_{1}$ in $Y_{2}$ is a projective line over $\mathbb{F}_{2}$, with self-intersection 1-1- 1. It is an exceptional divisor.
(IV.3.1) Suppose $\boxed{n_{1}\text{\ is odd}}$. Then $\lambda_{p_{2}}(Y_{2})=2$. After blowing-up $p_{2}$, the exceptional locus $\Theta_{3}$ is a union of two projective lines over $\mathbb{F}_{2}$, of multiplicity 222, intersecting each other at a rational point $p_{3}$. It intersects the strict transform of $\Theta_{2}$ at $p_{3}$. Moreover the point $p_{3}$ is regular in $Y_{3}$. The minimal regular model over $\mathbb{Z}_{2}$ of the genus 222 curve defined by Equation (3.2) has type $[2{\tt IV}-2]$ in [16], page 165.
(IV.3.2) Suppose $\boxed{n_{1}=2n_{2}\text{\ is even}}$, then $\lambda_{p_{2}}(Y_{2})=3$. We have an equation
$z_{2}^{2}+n_{2}s^{3}(1+x)^{3}z_{2}+x^{3}-ms^{4}(1+x)^{4}=0.$ (3.10)
with $p_{2}=(s=x=z_{2}=0)$. The blowing-up $Y_{3}\to Y_{2}$ along $p_{2}$ contains a unique singular point $p_{3}\in D_{+}(\tilde{s})$ defined by $s=x/s=z_{2}/s=0$, with multiplicity $\lambda_{p_{3}}(Y_{3})=4$. The exceptional locus $\Theta_{3}$ is isomorphic to $\mathbb{P}^{1}_{\mathbb{F}_{2}}$ and has multiplicity 444.
(IV.3.3) For the normalized blowing-up $Y_{4}\to Y_{3}$ along $p_{3}$, the equation of $Y_{4}$ over $D_{+}(\tilde{s})\subset Z_{4}$ is
$z_{3}^{2}+n_{2}s^{2}(1+sx_{1})^{3}z_{3}+x_{1}^{3}s-ms^{2}(1+sx_{1})^{4}=0.$
Blow-up-normalize $Y_{4}\to Y_{3}$ along $p_{3}$ gives an equation over $D_{+}(\tilde{s})$
$z_{4}^{2}+n_{2}(1+s^{2}x_{2})^{3}z_{3}+x_{2}^{3}-m(1+s^{2}x_{2})^{4}=0.$
If $\boxed{n_{2}\text{ \ is odd}}$ (so 444 divides exactly $n$), then the closed fiber of the Lipman’s resolution of the initial Weierstrass model is
where the last (elliptic) component is given by the equation
$w^{2}+w+x^{3}+\bar{m}=0$
over $\mathbb{F}_{2}$. After contracting the strict transforms of $\Theta_{1}$ and $\Theta_{3}$, we find the minimal regular model which has type $[2{\tt I}_{0-2}]$ in [16], page 159,
This type is denoted by $[2{\tt I}_{0}-3]$ in [11], pp. 71-72, for compatibility with the more general types $[2{\mathcal{K}}-\ell]$.
  • Remark 3.11
    The genus 222 curve in the above example is that of [11], Exemple 1 page 74. But the resolution of singularities there is incorrect. Nevertheless, the integer $d$ defined in op. cit., Théorème 1, page 70, is equal to 555 in Case (IV.3.3). It still satisfies the inequalities stated at the beginning of that example.
    To find the Artin conductor of the Jacobian of our genus 222 curve, we need not only the minimal resolution $X\to Y$, but also the integer $d$ mentioned above. The latter can be read from the closed fiber of the last $Z_{i}$ when there is no more singularities. We will come back to this question in [13].
3.6. Simultaneous resolution of singularities
Let $f:X\to Y$ be a finite morphism of integral noetherian excellent surfaces. S. Abhyankar [1] asked for the existence of resolutions of singularities $\tilde{X}\to X$ and $\tilde{Y}\to Y$ such that $f$ extends to a finite morphism $\tilde{X}\to\tilde{Y}$. He proved that the answer is no if $\deg f>3$, and is positive when $f$ is a Galois cover of degree 3$\leq$ 3≤ 3 of algebraic surfaces over a field of characteristic different from $\deg f$. For normal surfaces fibered over a discrete valuation ring, similar results are obtained in [14], §7 without assumption on the residue characteristics.
Corollary 3.12.
Let $f:X\to Y$ be a finite morphism of degree 222 of integral noetherian excellent surfaces. Then there are resolutions of singularities $\tilde{X}\to X$ and $\tilde{Y}\to Y$ such that $f$ extends to a finite morphism $\tilde{X}\to\tilde{Y}$.
Proof.
Replacing $Y$ with a resolution of singularities of $Y$ and then $X$ by the normalization of $Y$ in $K(X)$, we can assume that $Y$ is regular. Then $X\to Y$ is a normal double cover (Example 1). Applying Theorem 3.4 to $X\to Y$ then yields a simultaneous resolution of singularities. ∎
4. Variations of the multiplicities after normalized blowing-ups
In this section we will give a bound on the multiplicities of the points of $\tilde{Y}$ lying over $E$ in terms of $\lambda_{q_{0}}(Y)$ (Theorem 4.6). It says roughly that the sum of the new multiplicities is bounded by $\lambda_{q_{0}}(Y)$. This is done by comparing them to the multiplicities of the double cover $\tilde{Y}\times_{Z^{\prime}}E\to E$ when $\tilde{Y}\times_{Z^{\prime}}E$ is reduced, or of a “twisted” version of $\lambda_{q}$ otherwise.
4.1. Double covers of $\mathbb{P}^{1}_{k}$
We first give some properties of the multiplicities $\lambda_{q}$ on double covers of regular curves.
Proposition 4.1.
Let $E$ be an integral regular curve over a field $k$. Let $\Theta\to E$ be a double cover ($\Theta$ may be reducible) and let $\pi:\tilde{\Theta}\to\Theta$ be the normalization map.
  1. (1)
    Let $q\in E$ be a closed point $q\in E$. We have
    $\dim_{k}\left((\pi_{*}\mathcal{O}_{\tilde{\Theta}})_{q}/\mathcal{O}_{\Theta,q}\right)=[\lambda_{q}(\Theta)/2].$
  2. (2)
    Let $p\in\Theta$ be a point lying over $q$. If $\lambda_{q}(\Theta)$ is odd, then there is only one point $\tilde{p}\in\tilde{\Theta}_{p}$, and we have $k(p)=k(\tilde{p})$.
  3. (3)
    If $E$ is proper over $k$, then
    $p_{a}(\Theta)=p_{a}(\tilde{\Theta})+\sum_{q\in E}[\lambda_{q}(\Theta)/2][k(q):k],$
    where $p_{a}(C)=1-\chi(\mathcal{O}_{C})$ denotes the arithmetic genus of any proper variety $C$ over $k$.
Proof.
(1)-(2) Similar to [12], Lemme 6. The proof of (3) is similar to [9], Proposition 7.5.4. ∎
Next we consider the case $E=\mathbb{P}^{1}_{k}$. Fix an integer $r\geq 1$ and two homogeneous polynomials $a(s,t),b(s,t)\in k[s,t]$ such that
$\max\{2\deg a,\deg b\}=2r.$
Consider the curve $\Theta$ in the weighted projective space $\mathbb{P}(1,1,r)$ defined by the equation
$z^{2}+a(s,t)z+b(s,t)=0.$
Lemma 4.2.
Let $\Theta$ be as above and suppose that $\Theta$ is reduced. Denote by $\tilde{\Theta}\to\Theta$ the normalization map.
  1. (1)
    We have a finite flat morphism $\Theta\to\mathbb{P}^{1}_{k}$ of degree 222. Moreover $H^{0}(\Theta,\mathcal{O}_{\Theta})=k$ and $p_{a}(\Theta)=r-1$.
  2. (2)
    If $\Theta\to\mathbb{P}^{1}_{k}$ is generically étale then
    $\sum_{q\in\mathbb{P}^{1}_{k}}\lambda_{q}(\Theta)[k(q):k]\leq 2r.$
  3. (3)
    Suppose that $\Theta\to\mathbb{P}^{1}_{k}$ is purely inseparable.
    1. (a)
      If $H^{0}(\tilde{\Theta},\mathcal{O}_{\tilde{\Theta}})=k$, then
      $\sum_{q\in\mathbb{P}^{1}_{k}}2[\lambda_{q}(\Theta)/2][k(q):k]\leq 2r-2.$
    2. (b)
      Otherwise, we have
      $\sum_{q\in\mathbb{P}^{1}_{k}}2[\lambda_{q}(\Theta)/2][k(q):k]=2r,$
      and $\lambda_{q}(\Theta)$ is even if $k(q)=k$.
  4. (4)
    In all cases, $\lambda_{q}(\Theta)\leq 2r$.
Proof.
(1) The first part is clear and the second part can be proved by deforming the curve to a hyperelliptic curve of genus $r-1$, or by computing the Čech cohomology of $\mathcal{O}_{\Theta}$ with the affine covering $\{D_{+}(s),D_{+}(t)\}$.
(2) Denote by $\Delta(t)=\Delta_{\Theta/\mathbb{P}^{1}_{k}}\in k[t]$. For any closed point $q\in\operatorname{Spec}k[t]$, $\lambda_{q}(\Theta)\leq\operatorname{ord}_{q}(\Delta)$ by Proposition 2.6(2). We have
$\sum_{q\in\operatorname{Spec}k[t]}\lambda_{p}(\Theta)[k(q):k]\leq\deg\Delta(t)\leq 2r.$
We are done if $t=\infty$ corresponds to a regular point. Otherwise we have $\lambda_{\infty}(\Theta)\leq\deg t^{2r}\Delta(1/t)=2r-\deg\Delta(t)$.
(3.a) We have $p_{a}(\tilde{\Theta})=\dim H^{1}(\tilde{\Theta},\mathcal{O}_{\tilde{\Theta}})\geq 0$. The desired inequalities results from Proposition 4.1(3) and (1).
(3.b) Let $\tilde{k}=H^{0}(\tilde{\Theta},\mathcal{O}_{\tilde{\Theta}})$ with $k\subsetneq\tilde{k}$. Then $\tilde{\Theta}$ is a $\tilde{k}$-scheme. The morphism $\tilde{\Theta}\to\Theta\to\mathbb{P}^{1}_{k}$ induces a morphism $\tilde{\Theta}\to\mathbb{P}^{1}_{\tilde{k}}$ :
${\tilde{\Theta}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{P}^{1}_{\tilde{k}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Theta\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{P}^{1}_{k}.}$
We see that $\tilde{\Theta}\to\mathbb{P}^{1}_{\tilde{k}}$ has degree $2/[\tilde{k}:k]$, thus $[\tilde{k}:k]=2$, and $\tilde{\Theta}\to\mathbb{P}^{1}_{\tilde{k}}$ is birational hence an isomorphism. Therefore $p_{a}(\tilde{\Theta})=-1$ and
$\sum_{q\in\mathbb{P}^{1}_{k}}[\lambda_{q}(\Theta)/2][k(q):k]=r$
by Proposition 4.1 and (1). Note that $\tilde{k}/k$ is inseparable because $k(\Theta)=k(\tilde{\Theta})=\tilde{k}(t)$ is inseparable over $k(t)$.
Let $k(q)=k$ and suppose that $\lambda_{q}(\Theta)$ is odd. Let $p\in\Theta_{q}$ and let $\tilde{p}\in\tilde{\Theta}_{p}$. Then $k(p)=k(q)=k$ because $p$ is singular, and $k(\tilde{p})=k(p)=k$ by Proposition 4.1(2). But $\tilde{k}=H^{0}(\tilde{\Theta},\mathcal{O}_{\tilde{\Theta}})\subseteq k(\tilde{p})$, contradiction.
(4) is an immediate consequence of (2) and (3). ∎
4.2. Multiplicity along a regular closed subscheme
Let $\psi:Y\to Z$ be a normal double cover of an integral noetherian regular excellent scheme. Let $\Gamma$ be an integral regular closed subscheme of $Z$. We define in this subsection a multiplicity $\lambda_{\Gamma,q}(Y)$ which is just $\lambda_{q}(Y\times_{Z}\Gamma)$ if $Y\times_{Z}\Gamma\to\Gamma$ is a double cover (equivalently if $Y\times_{Z}\Gamma$ is reduced). If this condition is not satisfied, we have to “twist” $Y\times_{Z}\Gamma$ to get a double cover.
Lemma 4.3.
Let $A$ be integral noetherian regular and excellent. Let $B$ be a normal and free $A$-algebra of rank 222. Suppose that $A\to B$ has ramification index 222 above some prime ideal $\mathfrak{q}\subset A$ of height 111, and that $A/\mathfrak{q}$ is normal.
  1. (1)
    There exists a generator $y$ of $B$ such that $a\in\mathfrak{q}$, $b\in\mathfrak{q}\setminus\mathfrak{q}^{2}$.
  2. (2)
    Fix an $y$ given by (1). If $\mathfrak{m}\in V(\mathfrak{q})$ is a maximal ideal of $A$ such that $A_{\mathfrak{m}}/\mathfrak{q}A_{\mathfrak{m}}$ is regular, then $\lambda_{\mathfrak{m}}(B)=\lambda_{\mathfrak{m}}(y-c)$ for some $c\in\mathfrak{q}$ (depending in $\mathfrak{m}$).
Proof.
(1) Let $y$ be a generator of $B$. As $A\to B$ has ramification index 222 above $\mathfrak{q}$, $B/\mathfrak{q}B\simeq(A/\mathfrak{q})[T]/(T^{2}+\bar{a}T+\bar{b})$ is non-reduced. So $T^{2}+\bar{a}T+\bar{b}$ is a square in $\mathrm{Frac}(A/\mathfrak{q})[T]$. But $A/\mathfrak{q}$ is integrally closed, so $T^{2}+\bar{a}T+\bar{b}$ is in fact a square in $(A/\mathfrak{q})[T]$. Therefore translating $y$ by a suitable element of $A$ we have $a,b\in\mathfrak{q}$. Moreover, $b\notin\mathfrak{q}^{2}$ because $B\otimes_{A}A_{\mathfrak{q}}$ is normal.
(2) Suppose that for some $c_{1}\in A$ we have
$\lambda_{\mathfrak{m}}(y-c_{1})=\min\{2\operatorname{ord}_{\mathfrak{m}}(a+2c_{1}),\operatorname{ord}_{\mathfrak{m}}(b+(a+2c_{1})c_{1}-c_{1}^{2})\}>\ell:=\lambda_{\mathfrak{m}}(y).$
This implies that $\operatorname{ord}_{\mathfrak{m}}(c_{1})\geq\ell/2$ and $\operatorname{ord}_{\mathfrak{m}}(b-c_{1}^{2})>\ell$. Therefore
$c_{1}^{2}\in b+\mathfrak{m}^{\ell+1}\subset\mathfrak{q}+\mathfrak{m}^{\ell+1}.$
In $A/\mathfrak{p}$, we have $\bar{c}_{1}^{2}\in\bar{\mathfrak{m}}^{\ell+1}$. As $A/\mathfrak{p}$ is regular at $\bar{\mathfrak{m}}$, this implies that $\operatorname{ord}_{\bar{\mathfrak{m}}}(\bar{c}_{1})\geq r:=\lceil(\ell+1)/2\rceil$. In other words $c_{1}=c+s$ with $c\in\mathfrak{q}$ and $s\in\mathfrak{m}^{r}$. As $r>\ell/2$, we have $\lambda_{\mathfrak{m}}(y-c)=\lambda_{\mathfrak{m}}((y-c_{1})+s)>\ell$. Repeating this operation if necessary we will get a generator $y-c$ as desired, because $\lambda_{\mathfrak{m}}(B)$ is finite. ∎
  • Definition 4.4
    Let $\psi:Y\to Z$ be as in the beginning of this subsection. Let $\Gamma\subset Z$ be an integral regular closed subscheme of codimension 111 such that $\psi$ has ramification index 222 above the generic point of $\Gamma$. Then $\Theta:=Y\times_{Z}\Gamma\to\Gamma$ is finite flat of rank 222, but $\Theta$ is not reduced. So we can’t define the multiplicity $\lambda_{q}(\Theta)$ as in Definition 2.1. Instead we will define a “twisted” version for $\Theta$.
    We construct an effective divisor $D$ on $\Gamma$ as follows. Let $U\subseteq Z$ be an affine open subset such that $\mathcal{O}_{Z}(-\Gamma)|_{U}=s\mathcal{O}_{Z}(U)$ is free, and that $\psi^{-1}(U)\to U$ is free, defined by some equation
    $y^{2}+ay+b=0$
    such that $a,b\in s\mathcal{O}_{Z}(U)$ and $b\notin s^{2}\mathcal{O}_{Z}(U)$ (Lemma 4.3(1)). Note that any another similar equation of $\psi^{-1}(U)$ will give another $b^{\prime}=ub+s^{2}c$ with $u$ a unit and $c\in\mathcal{O}_{Z}(U)$. Then define $D|_{U\cap\Gamma}$ as the zero divisor of $(s^{-1}b)|_{U\cap\Gamma}$. It does not depend on the choice of the equation. This allows to define an effective divisor $D$ on $\Gamma$.
    Now for any $q\in\Gamma$ we put
    $\lambda_{\Gamma,q}(Y)=\operatorname{ord}_{q}(\mathcal{O}_{\Gamma}(-D)_{q})<+\infty$
    (order at $q$ of the ideal $\mathcal{O}_{\Gamma}(-D)_{q}\subseteq\mathcal{O}_{\Gamma,q}$). It will be crucial for the proof of Theorem 4.6(1) that we can compute $\lambda_{\Gamma,q}(Y)$ with the same equation for all points in a suitable open subset of $\Gamma$.
Lemma 4.5.
Let $Y\to Z$ and $\Gamma$ be as above. Let $\xi_{\Gamma}$ be the generic point of $\Gamma$, let $q\in\Gamma$ be a closed point and let $\Theta=Y\times_{Z}\Gamma$.
  1. (1)
    If $Y\to Z$ has ramification index 222 over $\xi_{\Gamma}$, then
    $\lambda_{\Gamma,q}(Y)\geq\lambda_{q}(Y)-1\geq 0.$
  2. (2)
    Suppose that $Y\to Z$ has ramification index 111 above $\xi_{\Gamma}$. Then $\Theta\to\Gamma$ is a double cover and $\lambda_{q}(\Theta)\geq\lambda_{q}(Y)$.
Proof.
We can suppose that $Z=\operatorname{Spec}A$ with $(A,\mathfrak{m})$ local. Observe that for any $\alpha\in A$,
$\operatorname{ord}_{\mathfrak{m}}(\alpha)\leq\operatorname{ord}_{\mathfrak{m}}(\alpha|_{\Gamma})$
where $\alpha|_{\Gamma}$ denotes the image of $\alpha$ in $\mathcal{O}(\Gamma)$.
(1) We can suppose that $Z=\operatorname{Spec}A$ is local. Let $s\in\mathfrak{m}$ be a generator of the ideal $\mathcal{O}_{Z}(-\Gamma)$. Let $y\in B$ be such that $\lambda_{\mathfrak{m}}(B)=\lambda_{\mathfrak{m}}(y)$ and $a,b\in sA$ (Lemma 4.3(2)). By definition
$\lambda_{q}(Y)\leq\operatorname{ord}_{q}(b)=\operatorname{ord}_{q}(b/s)+1\leq\operatorname{ord}_{q}((b/s)|_{\Gamma})+1=\lambda_{\Gamma,q}(Y)+1$
(the last equality comes from the definition of $\lambda_{\Gamma,q}(Y)$). On the other hands, as $a,b\in\mathfrak{m}$, we have $\lambda_{q}(Y)\geq 1$.
(2) Let $B=A[y]$ with $\lambda_{q}(Y)=\lambda_{\mathfrak{m}}(y)$. The projection $\Theta\to\Gamma$ is finite flat of rank 222. The scheme $\Theta$ is reduced because it is a local complete intersection (Lemma 2.2) and reduced at the generic points. Hence $\Theta\to\Gamma$ is a double cover. An equation of $\Theta$ is given by
$\bar{y}^{2}+\bar{a}\bar{y}+\bar{b}=0\mod s$
Then
$\lambda_{q}(\Theta)\geq\min\{2\operatorname{ord}_{q}(\bar{a}),\operatorname{ord}_{q}(\bar{b})\}\geq\min\{2\operatorname{ord}_{q}(a),\operatorname{ord}_{q}(b)\}=\lambda_{q}(Y)$
and the lemma is proved. ∎
4.3. Upper bound on the multiplicities
Theorem 4.6.
Let $Z$ be an integral noetherian excellent surface. Let $\psi:Y\to Z$ be a normal double cover. Let $p_{0}\in Y$ be a singular point, $q_{0}=\psi(p_{0})$.
Let $\phi:\tilde{Y}\to Z^{\prime}$ be as in Theorem 3.4 and let $E\subset Z^{\prime}$ be the exceptional divisor of $Z^{\prime}\to Z$. Let $\Theta:=\tilde{Y}\times_{Z^{\prime}}E$.
  1. (1)
    If $\lambda_{q_{0}}(Y)$ is odd, then $\lambda_{q}(\tilde{Y})\geq 1$ for all $q\in E$, and
    $\sum_{q\in E}(\lambda_{q}(\tilde{Y})-1)[k(q):k(q_{0})]\leq\sum_{q\in E}\lambda_{E,q}(\tilde{Y})[k(q):k(q_{0})]\leq\lambda_{q_{0}}(Y).$
    (See § 4.2 for the definition of $\lambda_{E,q}(\tilde{Y})$.)
  2. (2)
    Suppose that $\lambda_{q_{0}}(Y)$ is even. Then $\Theta\to E$ is a double cover.
    1. (a)
      If $\lambda_{q_{0}}(Y)=\operatorname{ord}_{q_{0}}(\Delta_{Y/Z})$, then
      $\sum_{q\in E}\lambda_{q}(\tilde{Y})[k(q):k(q_{0})]\leq\sum_{q\in E}\lambda_{q}(\Theta)[k(q):k(q_{0})]\leq\lambda_{q_{0}}(Y).$
    2. (b)
      Otherwise,
      $2\sum_{q\in E}[\lambda_{q}(\tilde{Y})/2][k(q):k(q_{0})]\leq 2\sum_{q\in E}[\lambda_{q}(\Theta)/2][k(q):k(q_{0})]\leq\lambda_{q_{0}}(Y).$
    3. (c)
      In both cases we have
      $\lambda_{q}(\tilde{Y})[k(q):k(q_{0})]\leq\lambda_{q}(\Theta)[k(q):k(q_{0})]\leq\lambda_{q_{0}}(Y)$
      for all $q\in E$.
Proof.
We can suppose that $Z=\operatorname{Spec}A$ is local. Let $\mathfrak{m}$ be the maximal ideal of $A$, $k=A/\mathfrak{m}$, let $\lambda_{\mathfrak{m}}(B)=\lambda_{\mathfrak{m}}(y)$ and $r=[\lambda_{\mathfrak{m}}(B)/2]$.
(1) The cover $\phi:\tilde{Y}\to Z^{\prime}$ has ramification index 222 at the generic point of $\Theta$ (Proposition 3.7(2)). Let $s,t$ be a system of generators of $\mathfrak{m}$. We use the notation of Proposition 3.1. By Lemma 4.3(1), we can suppose that $a,b\in sA$. The equation of $\tilde{Y}\times_{Z^{\prime}}D_{+}(\tilde{s})$ is
$y_{1}^{2}+a_{1}y_{1}+b_{1}=0,\quad a_{1}=a/s^{r},\ b_{1}=b/s^{2r}$
(Proposition 3.3). As $\operatorname{ord}_{\mathfrak{m}}(a)\geq r+1$, we have $a_{1}\in s\widetilde{A}_{(\tilde{s})}$. Write
$b=\sum_{0\leq i\leq 2r+1}\beta_{i}t^{i}s^{2r+1-i}+\epsilon_{2r+2},\quad\beta_{i}\in A,\ \epsilon_{2r+2}\in\mathfrak{m}^{2r+2}.$
Then, if $t_{1}=t/s$,
$b_{1}/s=\sum_{0\leq i\leq 2r+1}\beta_{i}t_{1}^{i}+sc,\quad c\in\widetilde{A}_{(\tilde{s})}$
Let $f(\bar{t}_{1})=\sum_{0\leq i\leq 2r+1}\bar{\beta}_{i}\bar{t}_{1}^{i}\in k[\bar{t}_{1}]$ be the class of $b_{1}/s$ mod $s$. By Lemma 4.5(1), for all $q\in E\cap D_{+}(\tilde{s})$, we have
$1\leq\lambda_{q}(\tilde{Y})\leq\lambda_{E,q}(\tilde{Y})=\operatorname{ord}_{q}(f(\bar{t}_{1})).$
Therefore
$\sum_{q\in E\cap D_{+}(\tilde{s})}(\lambda_{q}(\tilde{Y})-1)[k(q):k]\leq\deg f(\bar{t}_{1}).$
The only point missed is the rational point $\{s_{1}=0\}$ in $D_{+}(\tilde{t})$, where $s_{1}=s/t$. The same reasoning as above shows that
$\lambda_{s_{1}=0}(\tilde{Y})\leq\operatorname{ord}_{s_{1}=0}(\sum_{0\leq i\leq 2r+1}\bar{\beta}_{i}\bar{s}_{1}^{2r+1-i})=2r+1-\deg f(\bar{t_{1}}).$
This implies the desired inequality on the sum running through $E$.
(2) The fact that $\Theta\to E$ is a double cover is Lemma 4.5(2). Let $\tilde{A}=\oplus_{n\geq 0}\mathfrak{m}^{n}$. Then
$\tilde{A}\otimes_{A}k\simeq\mathrm{Gr}_{\mathfrak{m}}(A)\simeq k[s,t]$
because $A$ is regular. For any $f\in\mathfrak{m}^{n}$, if we consider $f$ as a homogeneous element $\tilde{f}\in\tilde{A}_{n}$, then the image $\tilde{f}(q_{0})$ of $\tilde{f}$ in $\tilde{A}\otimes_{A}k$ is 0 if $\operatorname{ord}_{\mathfrak{m}}(f)>n$, and is a homogeneous polynomial of degree $n$, equal to the initial form of $f$ otherwise. By Remark 3.3, $\Theta$ is the closed subscheme of $\mathbb{P}(1,1,r)$ defined by the equation
$w^{2}+\tilde{a}(q_{0})w+\tilde{b}(q_{0})=0$
where $a,b$ are respectively considered as elements of $\mathfrak{m}^{r}$ and $\mathfrak{m}^{2r}$. This is a double cover of $E$ of arithmetic genus $r-1$. So (2) is an immediate consequence of Lemmas 4.2 and 4.5. ∎
Corollary 4.7.
Keep the assumption of Theorem 3.4. For any $i\geq 1$, denote by $E_{i}\subseteq Z_{i}$ the exceptional divisor of $Z_{i}\to Z_{i-1}$. Then for all $q\in E_{i}$, we have
$\lceil\lambda_{q}(Y_{i})/2\rceil\leq\lceil\lambda_{q_{0}}(Y)/2\rceil.$
Proof.
By induction on $i$ and by using Theorem 4.6 (1) and (2.c). ∎
  • Example 4.8
    Let $Z=\operatorname{Spec}k[s,t]$ over a field $k$. Let $n\geq 3$. Consider $Y\subset\operatorname{Spec}k[s,t,y]$ defined by
    $y^{2}+t^{3}+s^{n}=0.$
    Let $q_{0}=(0,0)\in Z$. Then $\lambda_{q_{0}}(Y)=3$. We get a sequence of normalized blowing-ups (for $i+1<n/3$)
    $\to Y_{i+1}\to Y_{i}\to\dots\to Y_{1}\to Y_{0}=Y$
    where $Y_{i}$ has an affine open subset defined by
    $y_{i}^{2}+s^{n-3i+1}+st_{i}^{3}=0,\quad t_{i}=t/s^{i},y_{i}=y/s^{3m+1},$
    if $i=2m+1$, and by
    $y_{i}^{2}+s^{n-3i}+t_{i}^{3}=0,\quad t_{i}=t/s^{i},y_{i}=y/s^{3m},$
    if $i=2m$. The center of $Y_{i+1}\to Y_{i}$ is $p_{i}=\{y_{i}=s=t_{i}=0\}$ which has $\lambda_{p_{i}}(Y_{i})=4$ or 333 depending on whether $i$ is odd or even. There is no other singular points on $Y_{i}$.
4.4. The case when $\lambda_{p}(Y)=2$
Corollary 4.9.
Let $Y\to Z$ be as in Theorem 4.6. Let $q_{0}\in Z$ be a closed point with residue field $k$. Suppose that
$\lambda_{q_{0}}(Y)=2.$
Let $\Theta:=\tilde{Y}\times_{Z^{\prime}}E$.
  1. (1)
    The variety $\Theta$ is a reduced projective conic over $k$. There are two cases:
    1. (a)
      either $\Theta$ is regular (and integral). Then all points of $\Theta$ are regular in $\tilde{Y}$;
    2. (b)
      or $\Theta$ has a unique singular point $p_{1}$. Then $p_{1}$ is rational over $k$. There are two subcases.
      1. (i)
        $\Theta$ is union of two copies of $\mathbb{P}^{1}_{k}$ intersecting transversely at $p_{1}$;
      2. (ii)
        $\Theta$ is integral with normalization isomorphic to $\mathbb{P}^{1}_{\tilde{k}}$ for some quadratic extension $\tilde{k}/k$.
  2. (2)
    Suppose we are in Case (1.b) and that $p_{1}$ is singular in $Y_{1}$ (hence $\lambda_{p_{1}}(Y_{1})=2$). Let $Y_{2}\to Y_{1}=\tilde{Y}$ be the normalized blowing-up along $p_{1}$. Define similarly $\Theta_{2}\subset Y_{2}$.
    1. (a)
      The strict transform $\Theta^{\prime}$ of $\Theta$ in $Y_{2}$ is isomorphic to $\mathbb{P}^{1}_{k^{\prime}}$ for some reduced $k$-algebra $k^{\prime}$ of dimension 222, it intersects $\Theta_{2}$ transversely with intersection number $\Theta^{\prime}.\Theta_{2}=2$ (over $k$). Moreover, each irreducible component of $\Theta^{\prime}$ intersects an irreducible component of $\Theta_{2}$ and vice-versa.
    2. (b)
      In Case (1.b.i), $k^{\prime}\simeq k\times k$. Each irreducible component of $\Theta^{\prime}$ intersects exactly one irreducible component of $\Theta_{2}$ at a $k$-rational point. If $\Theta_{2}$ is irreducible, then it is isomorphic to $\mathbb{P}^{1}_{k}$.
    3. (c)
      In Case (1.b.ii), $k^{\prime}$ is a field, $\Theta_{2}$ is irreducible and $\Theta_{2}\cap\Theta^{\prime}$ consists in a $k^{\prime}$-rational point. Its normalization is isomorphic to $\mathbb{P}^{1}_{k^{\prime}}$ if it is singular.
  3. (3)
    The final resolution of singularity of $p_{0}$ has exceptional locus as follows:
    1. (a)
      In Case (1.b.i), all components are isomorphic to $\mathbb{P}^{1}_{k}$ and all intersection points are transverse and $k$-rational.
      (Read the resolution sequences from left to right.)
    2. (b)
      In Case (1.b.ii), all components but the last one are isomorphic to $\mathbb{P}^{1}_{k^{\prime}}$, all intersection points are transverse and $k^{\prime}$-rational. The last component is an irreducible conic (possibly singular).
Proof.
(1) is a consequence of Theorem 4.6.(2) and of the classification of projective conics over $k$. As $\Theta$ is a complete intersection in $Y$, if it is regular then it is contained in the regular locus of $Y$.
(2) By the above remark, $\Theta^{\prime}\cap\Theta$ is contained in the regular locus of $Y_{2}$. Parts (a) and (b) follow easily from Proposition 3.9.
(2.c) As $\Theta^{\prime}.\Theta_{2}$ has degree 222 over $k$ and $\Theta^{\prime}$ has no $k$-rational point, $\Theta^{\prime}\cap\Theta_{2}$ is just a $k^{\prime}$-rational point. By (a), $\Theta_{2}$ has only one irreducible component. Finally, if $\Theta_{2}$ is singular, its normalization is $\mathbb{P}^{1}$ over a quadratic extension of $k$. This extension is $k^{\prime}$ because $\Theta_{2}$ has a regular $k^{\prime}$-point.
(3) Apply inductively (2). ∎
4.5. Extended Tate’s algorithm
Suppose that $\lambda_{p_{0}}(Y)=3$. In Lipman’s resolution of singularities, the singular points lying over $p_{0}$ have multiplicities $\lambda\leq 4$ (Corollary 4.7). If for some reason these multiplicities are always bounded by 333, then one can repeat the same steps as in Tate’s algorithm [21] to get the complete resolution of $p_{0}$. The above condition on the multiplicities is satisfied if $Y$ is the minimal Weierstrass model of an elliptic curve over a discrete valuation ring as in [21].
If $Y$ is a minimal Weierstrass model of a genus 222 curve over a discrete valuation ring $R$, then $\lambda_{p_{0}}(Y)\leq 3$ if $p_{0}$ is rational over the residue field $k$ of $R$ ([10], Proposition 4.3.1), or $\lambda_{p_{0}}(Y)=4$ for a unique $p_{0}$ quadratic over $k$. So the singular points in Lipman’s sequence all have multiplicity bounded by 444 (Corollary 4.7). But even starting with $\lambda_{p_{0}}(Y)=3$ we can get a singular point with $\lambda=4$ (see Example 3.5, (IV)). The complete classification by Namikawa and Ueno [16] involves much more configurations then in the elliptic curves case.
5. Dualizing sheaves
We keep the hypothesis and notation of § 3.3. We study in this section the dualizing sheaves on the $Y_{i}$’s. Suppose further that the extension $K(Y)/K(Z)$ is separable and that we have a flat and finite type morphism $Z\to S$ to a noetherian regular scheme $S$.
Lemma 5.1.
Consider the commutative diagram
[…]
as in Theorem 3.4(2). Let $E$ be the exceptional divisor of $Z_{1}\to Z_{0}$ and $r=[\lambda_{p_{0}}(Y_{0})/2]$. Then we have canonical isomorphisms
$\omega_{Y_{1}/Z_{1}}\simeq\psi_{1}^{*}(\mathcal{O}_{Z_{1}}(-rE))\otimes c_{1}^{*}\omega_{Y_{0}/Z_{0}}$ (5.1)
and
$\omega_{Y_{1}/S}\simeq\psi_{1}^{*}(\mathcal{O}_{Z_{1}}(-(r-1)E))\otimes c_{1}^{*}\omega_{Y_{0}/S}$ (5.2)
Proof.
As $Y_{i}\to Z_{i}$ is a local complete intersection, $\omega_{Y_{i}/Z_{i}}$ is an invertible sheaf on $Y_{i}$. We have
$\omega_{Y_{1}/Z_{1}}\otimes(c_{1}^{*}\omega_{Y_{0}/Z_{0}})^{\vee}=\mathcal{O}_{Y_{1}}(D)$
for some Cartier divisor $D$ supported in the exceptional locus of $Y_{1}\to Y_{0}$. The determination of $D$ is local on $Z_{0}$. So we can suppose that $Z_{0}$ is affine and $Y_{0}$ is defined by an equation $y^{2}+ay+b=0$ with $\lambda_{p_{0}}(Y_{0})=\lambda_{p_{0}}(y)$. We then have
$\omega_{Y_{0}/Z_{0}}=\frac{1}{2y+a}\mathcal{O}_{Z_{0}}$
([9], Example 6.4.15). Consider the affine open subset $U_{1}$ of $Y_{1}$ defined by the equation
$y_{1}^{2}+a_{1}y+b_{1}=0$
with $y_{1}=y/s^{r},a_{1}=a/s^{r},b_{1}=b/s^{2r}$. Then $U_{1}=\psi_{1}^{-1}(V_{1})$ where $V_{1}=D_{+}(s)\subset Z_{1}$ (Proposition 3.3). We have
$\omega_{U_{1}/V_{1}}=\frac{1}{2y_{1}+a_{1}}\mathcal{O}_{U_{1}}=s^{r}\frac{1}{2y+a}\mathcal{O}_{U_{1}}=s^{r}c_{1}^{*}(\omega_{Y_{0}/Z_{0}})|_{U_{1}}.$
This implies that $D=\psi_{1}^{*}(-rE)$ and the first isomorphism is established.
The morphisms $Z_{0}\to S$ and $Y_{0}\to S$ are flat and local complete intersections ([9] Example 6.3.18 and Proposition 6.3.20(b)). The adjunction formula $\omega_{Y_{i}/S}=\psi_{0}^{*}(\omega_{Z_{i}/S})\otimes\omega_{Y_{i}/Z_{i}}$ ([9], Lemma 6.4.26(b)) and the canonical isomorphism $\omega_{Z_{1}/S}\simeq(b_{1}^{*}\omega_{Z_{0}/S})(E)$ ([9], Proposition 9.2.24) combined with the isomorphism (5.1) imply immediately the second isomorphism. ∎
Now consider a sequence as in Theorem 3.4(2)
[…]
For all $1\leq i\leq n$, denote by $p_{i-1}\in Y_{i-1}$ the center of $Y_{i}\to Y_{i-1}$, $E_{i}\subset Z_{i}$ the exceptional divisor of $Z_{i}\to Z_{i-1}$ as well as its strict transform in any $Z_{j}$, $j\geq i$. Define inductively a sequence of non-negative integers $(m_{i})_{1\leq i\leq n}$ as follows
  1. (1)
    $m_{1}=[\lambda_{p_{0}}(Y_{0})/2]-1$;
  2. (2)
    for all $1<i\leq n$,
    $m_{i}=([\lambda_{p_{i-1}}(Y_{i-1})/2]-1)+\sum_{0<j<i,\ E_{j}\ni p_{i-1}}m_{j}.$
  • Remark 5.2
    The index set in the above summation can be empty (if $p_{i-1}\in Y_{0}\setminus\{p_{0}\}$), or contains one $E_{j}$, or contains two components (if $p_{i-1}$ is an intersection point in the exceptional locus of $Y_{i}\to Y_{0}$).
Proposition 5.3.
Keep the assumptions of the beginning of this section and the above notation. Denote the composition
$\phi=\phi_{n}\colon\ Y_{n}\to Y_{n-1}\to\cdots\to Y_{0}$
and
$D_{n}=\sum_{1\leq i\leq n}m_{i}E_{i}.$
  1. (1)
    Then we have two canonical isomorphisms
    $\omega_{Y_{n}/S}\simeq\psi_{n}^{*}(\mathcal{O}_{Z_{n}}(-D_{n}))\otimes_{\mathcal{O}_{Y_{n}}}\phi^{*}(\omega_{Y_{0}/S}),$
    $\phi_{*}\omega_{Y_{n}/S}\simeq\phi_{*}(\psi_{n}^{*}(\mathcal{O}_{Z_{n}}(-D_{n})))\otimes_{\mathcal{O}_{Y_{0}}}\omega_{Y_{0}/S}.$
  2. (2)
    We have canonically
    $H^{0}(Y_{n},\omega_{Y_{n}/S})\subseteq H^{0}(Y_{n-1},\omega_{Y_{n-1}/S})\subseteq\cdots\subseteq H^{0}(Y_{0},\omega_{Y_{0}/S}).$
Proof.
(1) The first isomorphism is easily proved by induction on $n$. Indeed, the desired isomorphism is just Lemma 5.1 when $n=1$. To pass from $n$ to $n+1$, we just need to notice that the pull-back of $E_{j}\subseteq Z_{n}$ to $Z_{n+1}$ is $E_{j}$ if $p_{n}\notin E_{j}$, and is $E_{j}+E_{n}$ otherwise.
The second isomorphism is obtained by applying the projection formula to the first one, noting that $\omega_{Y_{0}/S}$ is invertible.
(2) It is enough to consider the case $n=1$. As $\psi_{1}$ is flat, we have
$\phi_{*}(\psi_{1}^{*}(\mathcal{O}_{Z_{1}}(-D_{1})))\subseteq\phi_{*}\mathcal{O}_{Y_{1}}=\mathcal{O}_{Y_{0}},$
the equality holds because $\phi$ is proper and birational, and $Y_{0}$ is normal. ∎
  • Remark 5.4
    Suppose that the generic fiber of $Y_{0}\to S$ is smooth.
    1. (1)
      The isomorphisms in the above proposition are equalities of subsheaves of the sheaf of rational differential forms $\Omega^{1}_{Y_{0}/S}\otimes_{\mathcal{O}_{Y_{0}}}K(Y_{0})$.
    2. (2)
      The global sections of $\omega_{Y_{n}/S}$ can be found as follows. For all $1\leq i\leq n$, denote by $\eta_{i}$ the generic point of $E_{i}$ and $\mathfrak{m}_{\eta_{i}}$ the maximal ideal of $\mathcal{O}_{Z_{n},\eta_{i}}$. Let $\omega\in H^{0}(Y_{0},\omega_{Y_{0}/S})$. Then $\omega\in H^{0}(Y_{n},\omega_{Y_{n}/S})$ if and only if for all singular points $p\in Y_{0}$ and for all $\eta_{i}$ lying over $\psi_{0}(p)\in Z_{0}$, we have
      $\omega\in\mathfrak{m}_{\eta_{i}}^{m_{i}}(\omega_{Y_{0}/S})_{p}.$
    This follows from the Hartogs type lemma ([9], Lemma 9.2.17(a)) and the second isomorphism of Proposition 5.3(1). We will come back to this question when $Z_{0}=\mathbb{P}^{1}_{S}$ in [13].
6. Desingularization algorithm
Let $A$ be a noetherian regular excellent domain of dimension 222. We will also assume $A$ is factorial (e.g. local). This assumption implies that locally free $A$-algebras $B$ of rank 222 are free (because the quotient $A$-module $B/A$ is locally free of rank 111 by Lemma 2.2). Let $B$ be an integral free $A$-algebra of rank 222. We denote as before $Z=\operatorname{Spec}A$ and $Y=\operatorname{Spec}B$.
6.1. Normalization
We determine the integral closure $B^{\prime}$ of $B$ when $\Delta_{B/A}\neq 0$. Write $B=A[y]$ with $y^{2}+ay+b=0$. Let $t$ be a prime divisor of $a^{2}-4b$. Denote by $\nu_{t}$ the normalized valuation on $\mathrm{Frac}(A)$ induces by $tA$.
As $A$ (hence $B$) is excellent, $B^{\prime}$ is finite over $B$. By Serre’s criterion $B^{\prime}$ is Cohen-Macaulay, hence finite flat of rank 222 over $A$. Therefore $B^{\prime}$ has a basis $\{1,z\}$ over $A$. There exist $u,v\in A$ such that
$y=uz+v.$ (6.1)
In particular $B=A[y]$ is normal at primes lying over $tA$ if and only if $t\nmid u$. Note that $u\mid a+2v,b+av+v^{2}$, hence $u\mid b-v^{2}$.
Normalization algorithm Let $y$ and $t$ be as above.
  1. (1)
    Let $r=[(\min\{2\nu_{t}(a),\nu_{t}(b)\})/2]$. Let $y_{1}=y/t^{r}$, $a_{1}=a/t^{r}$, $b_{1}=b/t^{2r}$. Then $a_{1},b_{1}\in A$ and
    $y_{1}^{2}+a_{1}y_{1}+b_{1}=0$
    with $B\subseteq B_{1}:=A[y_{1}]\subset L:=\mathrm{Frac}(B)$ and $\min\{2\nu_{t}(a_{1}),\nu_{t}(b_{1})\}\leq 1$.
  2. (2)
    Now replace $y$ with $y_{1}$, we can assume $\min\{2\nu_{t}(a),\nu_{t}(b)\}\leq 1$. Then $A[y]$ is normal at primes above $tA$ except possibly when $\nu_{t}(a)>0$, $\nu_{t}(b)=0$ and $\mathrm{char}(A/tA)=2$ (Lemma 3.2).
  3. (3)
    Assume that the above conditions hold. By the discussions following Equation (6.1), we see that $A[y]$ is not normal at some prime lying over $tA$ if and only if there exists $c\in A$ such that $t\mid b-c^{2}$ (equivalently $b$ is a square in $A/tA$).
  4. (4)
    Suppose that such a $c\in A$ exists. Let $y_{1}=y-c$. Then $A[y]=A[y_{1}]$ with
    $y_{1}^{2}+(a+2c)y_{1}+(b+ac+c^{2})=0$
    and $\nu_{t}(a+2c),\nu_{t}(b+ac+c^{2})>0$. Go back to Step (1) with this new equation.
When the above algorithm terminates (this always happens eventually because the valuation $\nu_{t}(u)$ of the factor $u$ in Equation (6.1) decreases strictly when we pass through Step (4) twice.), we go to another prime divisor of $a^{2}-4b$. At the end we get $B^{\prime}$ together with a basis $\{1,z\}$.
6.2. Finding the singular points
We are given a normal $B=A[y]$ with
$y^{2}+ay+b=0.$
Singular points of $Y$. Even though there are only finitely many singular points, there is no obvious way to find them for general $A$.
- If $A$ (hence $B$) is a finite type algebra over some perfect field, we can use Jacobian criterion.
- Suppose that $Z$ is smooth with integral fibers over some Dedekind domain $R$ (e.g. $A=R[x]$) and that $B\otimes_{A}\mathrm{Frac}(R)$ is regular. We can look for singular points in the closed fibers of $Y\to\operatorname{Spec}R$ (which are curves over fields). The singular points of $Y$ are among them.
However, if $B\otimes_{R}{k(w)}$ is non-reduced for some closed point $w\in\operatorname{Spec}R$, all points of the fiber $Y_{w}$ are singular. The worst situation is when $\Delta_{B/A}(w)=0$. Suppose $R$ is local of dimension 111 with a uniformizing element $\pi$. Then $\pi\mid a$ and $b$ is a square in $A/\pi A$. Let $b\equiv c^{2}\mod\pi$. Replace $y$ with $y+c$, then we get an equation with $\pi\mid a,b$. As $B$ is normal, $\pi^{2}\nmid b$. The singular points of $Y$ are exactly those lying over the zeros of $b/\pi$ in the curve $\operatorname{Spec}(A\otimes_{R}k(w))$.
Singular points in the normalized blowing-ups. Let $q_{0}\in Z$ be such that $\lambda_{q_{0}}(Y)\geq 2$. Let $k=k(q_{0})$.
Let $\tilde{Y}\to Y$ be the normalized blowing-up at the point $p_{0}$ lying over $q_{0}$ (see § 3). We search for singular points of $\tilde{Y}$ lying over the exceptional locus $E\subset Z^{\prime}\stackrel{{\scriptstyle\mathrm{bl}_{\mathfrak{m}}}}{{\longrightarrow}}Z$. Let $a_{1},b_{1}\in\widetilde{A}_{(\tilde{s})}$ be as in Proposition 3.3. We first look for singular points lying over $D_{+}(\tilde{s})$. There are two cases.
[…]
(2) $s\mid a_{1}^{2}+4b_{1}$. Then we have $s\mid a_{1}$ (we take $a=0$ if $\mathrm{char}(k)\neq 2$).
(2.1) If $s\mid b_{1}$ then $s^{2}\nmid b_{1}$. The singular points correspond exactly to the zeros of $\overline{s^{-1}b_{1}}\in\widetilde{A}_{(\tilde{s})}/(s)\simeq k[t_{1}]$.
(2.2) If $\mathrm{char}(k)=2$, and $\bar{b}_{1}\in\widetilde{A}_{(\tilde{s})}/(s)$ is a square. Translate $y_{1}$ by an element of $\widetilde{A}_{(\tilde{s})}$ to make $s\mid b_{1}$. Go back to (2.1).
[…]
(3) It remains the points lying over $V_{+}(\tilde{s})\cap E$. Let $t\in\mathfrak{m}$ be such that $s,t$ generate $\mathfrak{m}A_{\mathfrak{m}}$. Then $V_{+}(\tilde{s})\cap E=V_{+}(\tilde{s})\cap D_{+}(\tilde{t})\cap E$ is the point $s_{1}=t=0$ in $D_{+}(\tilde{t})$ (where $s_{1}=s/t$). Check similarly whether it induces a singular point of $\tilde{Y}$
6.3. Desingularization algorithm
Let $Z$ be an integral noetherian excellent regular surface. Let $Y\to Z$ be a double cover. We work locally on $Z$ and suppose $Z=\operatorname{Spec}A$ is affine.
(1) Find the normalization of $Y$ (§ 6.1), then suppose $Y$ is normal.
(2) Find the singular points of $Y$ (§ 6.2).
(3) Fix a singular point lying over a closed point $q_{0}\in Z$. Compute $\lambda_{q_{0}}(Y)$ using Lemma 2.10.
(4) Consider the normalized blowing-up $\tilde{Y}\to Y$ along the point $p_{0}$ lying over $q_{0}$. Use the affine equation of $\tilde{Y}$ given Proposition 3.3 to find the singular points of $\tilde{Y}$ lying over $p_{0}$.
(5) Repeat the above process for all singular points of $Y$. At the end we get a new normal double cover of an integral noetherian excellent regular surface.
(6) Go back to (2).
6.4. Presentation of a regular ring and of its Rees algebras
The most computationally complicated part of our algorithm should be the determination of $\operatorname{ord}_{\mathfrak{m}}$ and of the initial form of an element of $A$ (Lemma 2.10(2.b)). For these purposes we need an as simple as possible presentation (by generators and relations) of $A$ and of the rings appearing after blowing-up $\operatorname{Spec}A$ along maximal ideals.
We will restrict ourselves to the case when $A$ is of finite type over some principal ideal domain $R$ (e.g. $\mathbb{Z}$, $\mathbb{F}_{p}[T]$). Write
$A=R[U_{1},\dots,U_{n}]/I$
for some $n\geq 2$. Let $\mathfrak{m}\subset A$ be a maximal ideal of height 222 generated by $s,t$. Then
$\widetilde{A}_{(\tilde{s})}=A[T_{1}]/(sT_{1}-t)=R[U_{1},\dots,U_{n},T_{1}]/(I,ST_{1}-T)$
where $S,T\in R[U_{1},\dots,U_{n}]$ are liftings of $s,t$ respectively. Thus in general we need to add a new variable and a new relation to represent the $R$-algebra $\tilde{A}_{(\tilde{s})}$. However this is not necessary in some particular cases.
Suppose that we can write
$A=A_{0}[S,T]/I$
for some $R$-algebra $A_{0}$ such that $\mathfrak{m}$ is generated by the classes $s,t$ of $S,T$ in $A$. Then
$\widetilde{A}_{(\tilde{s})}=A_{0}[S,T,T_{1}]/(I,ST_{1}-T)=A_{0}[S,T_{1}]/I^{\prime}$
where $I^{\prime}$ is obtained by replacing $T$ with $ST_{1}$ in the polynomials in $I$. Similarly
$\tilde{A}_{(\tilde{t})}=A_{0}[S,T,S_{1}]/(I,TS_{1}-S)=A_{0}[T,S_{1}]/I^{\prime\prime}$
where $I^{\prime\prime}$ is obtained by replacing $S$ with $TS_{1}$ in the polynomials in $I$.
Let $\mathfrak{m}_{q}$ be a maximal ideal of $\widetilde{A}_{(\tilde{s})}$ containing $s$. It corresponds to a closed point $q$ of $E\cap D_{+}(\tilde{s})$. Suppose that $q$ is rational over $k:=A/\mathfrak{m}$. Then there exists $\theta\in A_{0}$ such that $\mathfrak{m}_{q}$ is generated by the classes of $S,T_{1}-\theta$. After translating $T_{1}$ by $\theta$, we can suppose $\theta=0$. Then the blowing-up of $\operatorname{Spec}\widetilde{A}_{(\tilde{s})}$ along $q$ is covered by the spectra of
$A_{0}[S,T_{2}]/J,\quad A_{0}[T_{1},S_{2}]/J^{\prime}$
where $J$ is obtained by substituting $ST_{2}$ to $T_{1}$ in the polynomials of $I^{\prime}$, and $J^{\prime}$ is obtained by substituting $T_{1}S_{2}$ to $S$ in the polynomials of $I^{\prime}$.
Corollary 6.1.
Let $R$ be a PID. Suppose $A=R[x]$. Let $\pi\in R$ be an irreducible element and let $k=R/(\pi)$. If we blow-up $\operatorname{Spec}A$ along a rational point $q$ of $\operatorname{Spec}k[x]$, and then successively blow-up $k$-rational points in the exceptional divisors, then all intermediate schemes are covered by regular affine schemes of the form
$\operatorname{Spec}(R[S,T]/(f(S,T)-\pi))$
with $f(S,T)\in(S,T)R[S,T]$.
Proof.
We can suppose that $q$ corresponds to the maximal ideal $(x,\pi)\subset R[x]$. Let $A_{0}=R$. We have $A=A_{0}[S,T]/(f(S,T)-\pi)$ with $f(S,T)=T$. Then apply the above discussions. ∎